### v1

# Poisson Midpoint Method for Log Concave Sampling: Beyond the Strong Error Lower Bounds

Abstract
We study the problem of sampling from strongly log-concave distributions over $\mathbb{R}^{d}$ using the Poisson midpoint discretization (a variant of the randomized midpoint method) for overdamped/underdamped Langevin dynamics. We prove its convergence in the 2-Wasserstein distance ($W_{2}$), achieving a cubic speedup in dependence on the target accuracy ($\epsilon$) over the Euler-Maruyama discretization, surpassing existing bounds for randomized midpoint methods. Notably, in the case of underdamped Langevin dynamics, we demonstrate the complexity of $W_{2}$ convergence is much smaller than the complexity lower bounds for convergence in $L^{2}$ strong error established in the literature.
1 Introduction
Sampling from a density $\pi(x)\propto\exp(-F(x))$ over $\mathbb{R}^{d}$ is of fundamental interest in physics, economics, and finance johannes2010mcmc ; von2011bayesian ; kobyzev2020normalizing . Applications in computer science include volume computation vempala2010recent and bandit optimization russo2018tutorial . A popular approach is Langevin Monte Carlo (LMC) which is the Euler-Maruyama discretization of the continuous time Ito Stochastic Differential Equation (SDE) called (overdamped/underdamped) Langevin Dynamics. The convergence of LMC has been extensively studied in the literature durmus2018analysislangevinmontecarlo ; vempala2022rapidconvergenceunadjustedlangevin ; erdogdu2021convergencelangevinmontecarlo ; cheng2017convergencelangevinmcmckldivergence ; cheng2018-underdamped-lmc ; dalalyan2020sampling under various assumptions on the potential function $F$ such as log concavity and isoperimetry.
The randomized midpoint discretization for Langevin dynamics (RLMC), introduced by shen2019randomizedmidpointmethodlogconcave and developed further in yu2023langevinmontecarlostrongly ; he2021ergodicitybiasasymptoticnormality ; altschuler2024shiftedcompositioniiilocal considers a more sophisticated alternative to Euler-Maruyama discretization, reducing the bias while introducing variance, leading to faster convergence bounds than for LMC. The work of Cao et. al Cao_2021 established lower bounds for the computational complexity for convergence in the strong $L^{2}$ error for a class of sampling algorithms, which includes RLMC and LMC. This demonstrated that the computational complexity of RLMC with underdamped Langevin dynamics is optimal in dimension and accuracy (up to log factors) in terms of the strong $L^{2}$ error.
Strong $L^{2}$ error is the $L^{2}$ distance between the continuous time Ito SDE solution at time $T$ and the sampling algorithm output whenever they are driven by same Brownian motion. However, sampling algorithm guarantees generally consider ‘weak’ notions of distance such as total variation distance, Wasserstein distance, or the KL divergence between the law of algorithm output and the target. In particular, Wasserstein-2 distance bounds can consider the $L^{2}$ distance between algorithm output and the continuous time SDE driven by different but arbitrarily coupled Brownian motions.
The Poisson Midpoint Method for Langevin dynamics (PLMC) was introduced in kandasamy2024poissonmidpointmethodlangevin as a variant of RLMC. While kandasamy2024poissonmidpointmethodlangevin considered the convergence of PLMC under general conditions (beyond strong log-concavity) for the total variation distance via entropic central limit theorem style arguments, we consider a sharp analysis via coupling arguments for the case of strong log-concave sampling to obtain better convergence guarantees. This utilizes the sharp bound on the Wasserstein distance between a Gaussian random-variable and a perturbed Gaussian random variable used in Alex Zhai’s proof of the Central Limit Theorem in the Wasserstein distance zhai2017highdimensionalcltmathcalw2distance .
This allows us to obtain guarantees for convergence in the 2-Wasserstein distance which obtain a cubic improvement in the accuracy $\epsilon$ over the Euler-Maruyama discretization. This is provably better than the established lower bounds for the strong $L^{2}$ error in Cao_2021 , for this class of algorithms. This demonstrates that newer techniques are necessary to understand the limits of sampling.
2 Notation and problem setup
[…]
Given a sequence of probability measures $\mu_{i}$ over $\mathcal{X}_{i}$, for $i\in[n]$, a coupling is a probability measure $\Gamma$ over the product space $\prod_{i}\mathcal{X}_{i}$ such that the marginal over $\mathcal{X}_{j}$ is $\mu_{j}$. A sequence of random variables $(X_{i}\sim\mu_{i})$ are coupled if they are defined over a common probability space, since their joint law is a coupling of $(\mu_{i})_{i\in[n]}$. The Wasserstein-2 distance between $\mu$ and $\nu$ is given by
[…]
where $\Gamma(\mu,\nu)$ denotes the set of couplings of $\mu$ and $\nu$. We make the following assumptions on $F$.
Assumption 1.
The function $F:\mathbb{R}^{d}\to\mathbb{R}$ is $\alpha$ strongly convex and $L$ smooth for some $\alpha,L>0$. That is, $F$ is twice continuously differentiable over $\mathbb{R}^{d}$ and for every $x,y\in\mathbb{R}^{d}$, we have: $F(y)-F(x)\geq\langle\nabla F(x),y-x\rangle+\frac{\alpha}{2}\|x-y\|^{2}$ and $\|\nabla F(x)-\nabla F(y)\|\leq L\|x-y\|$.
The target distribution, given by the density $\pi(x)\propto\exp(-F(x))$, is then called strongly log-concave. Our goal is to sample a random variable $X\sim\mu$ such that:111scaling $\frac{d}{\alpha}$ as considered in shen2019randomizedmidpointmethodlogconcave .
$W_{2}^{2}(\mu,\pi)\leq\frac{\epsilon^{2}d}{\alpha}.$ (1)
We define the condition number $\kappa:=\frac{L}{\mu}$. Our notion of complexity is the number of gradient calls of $F$, in terms of the problem parameters $\kappa,d$ and $\epsilon.$ The remainder of this section closely follows the discussion in kandasamy2024poissonmidpointmethodlangevin , since we have the same problem setup.
2.1 Langevin Monte-Carlo
Consider a general, discrete time stochastic process $(X_{t})_{t\in\mathbb{N}}$ in $\mathbb{R}^{d}$, with step-size $\eta$.
$X_{t+1}=A_{\eta}X_{t}+G_{\eta}b(X_{t},t\eta)+\Gamma_{\eta}Z_{t}.$ (2)
Here $Z_{t}$ are independent standard Gaussians, $b(x,t):\mathbb{R}^{d}\times\mathbb{R}^{+}\to\mathbb{R}^{d}$ is a drift function; $A_{\eta}$ and $G_{\eta}$ are $d\times d$ matrices. We focus on two specific cases of the above.
Overdamped LMC: Here $A=\mathbf{I}_{d}$, $G_{\eta}=\eta\mathbf{I}_{d}$, $\Gamma_{\eta}=\sqrt{2\eta}\mathbf{I}_{d},$ and $b(x,t)=-\nabla F(x)$ for some potential $F.$ This is the Euler-Maruyama discretization of Overdamped Langevin dynamics:
$dX_{t}=-\nabla F(X_{t})dt+\sqrt{2}dB_{t}.$
This SDE converges to the stationary distribution $\pi(x)\propto\exp(-F(x))$ roberts1996exponential .
Underdamped LMC: Here $X_{t}=[U_{t},V_{t}]^{T}\in\mathbb{R}^{2d}$, with $U_{t}$ and $V_{t}$ denoting position and momentum respectively. The other terms in the recurrence are defined as follows.
$A_{\eta}=\begin{bmatrix}\mathbf{I}_{d}&\tfrac{1}{\gamma}(1-e^{-\gamma\eta})\mathbf{I}_{d}\\ 0&e^{-\gamma\eta}\mathbf{I}_{d}\end{bmatrix},\text{ }G_{\eta}=\begin{bmatrix}\tfrac{1}{\gamma}(\eta-\tfrac{1}{\gamma}(1-e^{-\gamma\eta}))\mathbf{I}_{d}&0\\ \tfrac{1}{\gamma}(1-e^{-\gamma\eta})\mathbf{I}_{d}&0\end{bmatrix},\text{ }b(X_{t})=\begin{bmatrix}-\nabla F(U_{t})\\ 0\end{bmatrix},$
$\Gamma_{\eta}^{2}:=\begin{bmatrix}\frac{2}{\gamma}\left(\eta-\frac{2}{\gamma}(1-e^{-\gamma\eta})+\frac{1}{2\gamma}(1-e^{-2\gamma\eta})\right)\mathbf{I}_{d}&\tfrac{1}{\gamma}(1-2e^{-\gamma\eta}+e^{-2\gamma\eta})\mathbf{I}_{d}\\ \tfrac{1}{\gamma}(1-2e^{-\gamma\eta}+e^{-2\gamma\eta})\mathbf{I}_{d}&(1-e^{-2\gamma\eta})\mathbf{I}_{d}\end{bmatrix}.$
This is the Euler-Maruyama discretization of the underdamped Langevin dynamics:
$dU_{t}=V_{t}dt,\text{ }dV_{t}=-\gamma V_{t}-\nabla F(U_{t})dt+\sqrt{2}dB_{t}$ (3)
The stationary distribution of these dynamics is $\pi(u,v)\propto\exp(-F(u)-\frac{1}{2}||v||^{2}).$ eberle2019couplings ; dalalyan2020sampling
2.2 The Poisson midpoint method
For $K\in\mathbb{N}$. Consider the dynamics in (2) with step size $\eta/K$, $X_{t}$. The Poisson midpoint method uses a randomized algorithm to approximate $K$ steps of the dynamics in with just 1 step of step size $\eta$ using 2 evaluations of the drift $b()$ in expectation. Denoting the iterates of the Poisson midpoint method as $\tilde{X}_{tK}$, this means $\tilde{X}_{tK}$ approximates $X_{tK}$. The work of kandasamy2024poissonmidpointmethodlangevin gives an efficient implementation of PLMC, with 2 evaluations of the drift $b()$ per $K$ steps to obtain $\tilde{X}_{(t+1)K}$ directly from $\tilde{X}_{tK}$. We consider the ‘fictitious’ interpolation $\hat{\tilde{X}}_{tK+i}$ (which approximates $X_{tK+i}$) to enable mathematical analysis. The iterates run in batches of size $K,$ denoted ${\tilde{X}}_{tK+i}$ with $0\leq i<K$. Let $Y_{tK+i}$ denote independent standard Gaussians. Given $\tilde{X}_{tK}$, the 0-th iterate of a batch, we define the cheap interpolation
$\hat{\tilde{X}}_{tK+i}=A_{\frac{\eta i}{K}}\tilde{X}_{tK}+G_{\frac{\eta i}{K}}b(\tilde{X}_{tK},t\eta)+\sum_{j=0}^{i-1}A_{\frac{\eta(i-j-1)}{K}}\Gamma_{\tfrac{\eta}{K}}Y_{tK+j}.$
Let $H_{i,t}$ be a sequence of independent Bernoulli random variables, with parameter $1/K.$ The Poisson midpoint iterates are then defined as follows, for $0\leq i<K$.
$\tilde{X}_{tK+i+1}=A_{\tfrac{\eta}{K}}\tilde{X}_{tK+i}+G_{\tfrac{\eta}{K}}\left[b(\tilde{X}_{tK},t\eta)+KH_{t,i}(b(\hat{\tilde{X}}_{tK+i},\tfrac{(tK+i)\eta}{K})-b(\tilde{X}_{tK},t\eta))\right]+\Gamma_{\tfrac{\eta}{K}}Y_{tK+i}.$
Define $\tilde{Z}_{tK+i}=\Gamma_{\frac{\eta}{K}}^{-1}G_{\tfrac{\eta}{K}}\big{[}b(\tilde{X}_{tK},t\eta)-b({\tilde{X}}_{tK+i},t\eta)+KH_{t,i}(b(\hat{\tilde{X}}_{tK+i},\tfrac{(tK+i)\eta}{K})-b(\tilde{X}_{tK},t\eta))\big{]},$ so that the recursion can be written as
$\tilde{X}_{tK+i+1}=A_{\frac{\eta}{K}}{\tilde{X}}_{tK+i}+G_{\frac{\eta}{K}}b({\tilde{X}}_{tK+i},t\eta)+\Gamma_{\frac{\eta}{K}}\tilde{Z}_{tK+i}.$
We consider this approximation in the two specific cases of Langevin Dynamics to obtain Poisson Langevin Monte-Carlo (PLMC).
Overdamped PLMC: Let $Z_{tK+i}$ be independent standard Gaussians, and $X_{tK+i}$ denote the iterates of Langevin monte-carlo, with step-size $\eta/K.$
$X_{tK+i+1}=X_{tK+i}-\tfrac{\eta}{K}\nabla F(X_{tK+i})+\sqrt{\tfrac{2\eta}{K}}Z_{tK+i}.$
Let $Y_{tK+i}$ be independent standard Gaussians. PLMC is defined via the following recurrences.
$\hat{\tilde{X}}_{tK+i}=\tilde{X}_{tK}-\tfrac{\eta i}{K}\nabla F(\tilde{X}_{tK})+\sqrt{\frac{2\eta}{K}}\sum_{j=0}^{i}Y_{tK+j}$
$\tilde{X}_{tK+i+1}=\tilde{X}_{tK+i}+\eta H_{i,t}(\nabla F(\tilde{X}_{tk})-\nabla F(\hat{\tilde{X}}_{tK+i}))-\tfrac{\eta}{K}\nabla F(\tilde{X}_{tK})+\sqrt{\tfrac{2\eta}{K}}Y_{tK+i}$
Underdamped PLMC: Let $Z_{tK+i}$ be independent standard Gaussians in $\mathbb{R}^{2d},$ and $X_{tK+i}=[U_{tK+i},V_{tK+i}]^{T}$ denote the iterates of underdamped Langevin Monte-Carlo with step-size $\eta/K$.
$X_{tK+i+1}=A_{\frac{\eta}{K}}{\tilde{X}}_{tK+i}+G_{\frac{\eta}{K}}b({\tilde{X}}_{tK+i})+\Gamma_{\frac{\eta}{K}}Z_{tK+i},$
$A_{h}:=\begin{bmatrix}\mathbf{I}_{d}&\tfrac{1}{\gamma}(1-e^{-\gamma h})\mathbf{I}_{d}\\ 0&e^{-\gamma h}\mathbf{I}_{d}\end{bmatrix},\text{ }G_{h}:=\begin{bmatrix}\tfrac{1}{\gamma}(h-\tfrac{1}{\gamma}(1-e^{-\gamma h}))\mathbf{I}_{d}&0\\ \tfrac{1}{\gamma}(1-e^{-\gamma h})\mathbf{I}_{d}&0\end{bmatrix},\text{ }b(X)=\begin{bmatrix}-\nabla F(U)\\ 0\end{bmatrix}$
$\Gamma_{h}^{2}:=\begin{bmatrix}\frac{2}{\gamma}\left(h-\frac{2}{\gamma}(1-e^{-\gamma h})+\frac{1}{2\gamma}(1-e^{-2\gamma h})\right)\mathbf{I}_{d}&\tfrac{1}{\gamma}(1-2e^{-\gamma h}+e^{-2\gamma h})\mathbf{I}_{d}\\ \tfrac{1}{\gamma}(1-2e^{-\gamma h}+e^{-2\gamma h})\mathbf{I}_{d}&(1-e^{-2\gamma h})\mathbf{I}_{d}\end{bmatrix},$
Let $Y_{tK+i}$ be independent standard Gaussians in $\mathbb{R}^{2d}$. Underdamped PLMC is defined by the following recurrences.
$\tilde{X}_{tK+i+1}=A_{\frac{\eta}{K}}{\tilde{X}}_{tK+i}+G_{\frac{\eta}{K}}b({\tilde{X}}_{tK+i})+\Gamma_{\frac{\eta}{K}}\tilde{Z}_{tK+i},$
$\tilde{Z}_{tK+i}=Y_{tK+i}+\Gamma_{\frac{\eta}{K}}^{-1}G_{\frac{\eta}{K}}[b(\hat{\tilde{X}}_{tK+i})-b({\tilde{X}}_{tK+i})]+K(H_{i,t}-1/K)\Gamma_{\frac{\eta}{K}}^{-1}G_{\frac{\eta}{K}}[b(\hat{\tilde{X}}_{tK+i})-b(\tilde{X}_{tK})],$
and $\hat{\tilde{X}}_{tK+i}=A_{\frac{i\eta}{K}}\tilde{X}_{tK}+G_{\frac{i\eta}{K}}b(\tilde{X}_{tK})+\sum_{j=0}^{i-1}A_{\frac{\eta(i-j-1)}{K}}\Gamma_{\tfrac{\eta}{K}}Y_{tK+j}.$
2.3 Prior work
Table 1: Convergence bounds for discretized Langevin Dynamics. In case of LSI, $\kappa=L\times$ LSI constant. The scaling of $W_{2}^{2}$ is different from (1) to compare with TV and KL bounds, and $p\in\mathbb{N}$ is arbitrary.
Overdamped Langevin Dynamics
Algorithm Assumption Metric Steps
LMC (durmus2018analysislangevinmontecarlo ) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa d}{\epsilon^{2}}$
RLMC (shen2019randomizedmidpointmethodlogconcave ; yu2023langevinmontecarlostrongly ) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa\sqrt{d}}{\epsilon}+\frac{\kappa^{4/3}d^{1/3}}{\epsilon^{2/3}}$
RLMC (altschuler2024shiftedcompositioniiilocal ) Log-Concave $\mathsf{KL}\leq\epsilon^{2}$ $\frac{\kappa\sqrt{d}}{\epsilon}$
RLMC (altschuler2024shiftedcompositioniiilocal ) LSI $\mathsf{KL}\leq\epsilon^{2}$ $\frac{\kappa^{3/2}\sqrt{d}}{\epsilon}$
PLMC (Ours) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa^{4/3}d^{1/3}+\kappa d^{2/3}}{\epsilon^{2/3}}$
Underdamped Langevin Dynamics
Algorithm Assumption Metric Steps
LMC (dalalyan2020sampling ) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa^{3/2}\sqrt{d}}{\epsilon}$
RLMC (shen2019randomizedmidpointmethodlogconcave ; yu2023langevinmontecarlostrongly ) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa d^{1/3}}{\epsilon^{2/3}}+\frac{\kappa^{7/6}d^{1/6}}{\epsilon^{1/3}}$
PLMC (kandasamy2024poissonmidpointmethodlangevin ) LSI $\mathsf{TV}\leq\epsilon$ $\frac{\kappa^{\frac{17}{12}}d^{\frac{5}{12}}}{\sqrt{\epsilon}}$
PLMC (Ours) Log-Concave $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ $\frac{\kappa^{7/6}d^{1/3}}{\epsilon^{1/3}}+\frac{\kappa^{\frac{11p+6}{8p+6}}d^{\frac{3p+2}{8p+6}}}{\epsilon^{\frac{p+2}{4p+3}}}$
Recent works have focused on the rigorous theoretical analysis of classical and popular MCMC algorithms to establish complexity bounds and theoretical limits. The prototypical case of Overdamped LMC has been studied when the target $\pi$ is strongly log-concave and more generally $\pi$ when satisfies isoperimetric inequalities dalalyan2017theoretical ; durmus2017unadjusted ; durmus2018analysislangevinmontecarlo ; vempala2022rapidconvergenceunadjustedlangevin ; erdogdu2021-tail-growth ; mou2022improved ; sinho-non-log-concave-lmc . Underdamped LMC has been considered as a faster alternative. This case too has been well studied when $\pi$ is strongly log-concave and when $\pi$ satisfies isoperimetric inequalities cheng2018-underdamped-lmc ; dalalyan2020sampling ; ganesh2020faster ; ma2021there ; zhang2023improveddiscretizationanalysisunderdamped
LMC is the Euler-Maruyama discretization of continuous time Langevin Dynamics, which can lead to sub-optimal convergence due to statistical bias in the approximation. Thus, shen2019randomizedmidpointmethodlogconcave introduced the randomized midpoint method for LMC (RLMC) which reduces the bias in the approximation by introducing a randomized estimator at the cost of higher variance. RLMC does not involve higher order derivatives of $\nabla F$ as in Runge-Kutta schemes for SDEs kloeden1992stochastic - allowing its use for generative modeling with denoising diffusion models kandasamy2024poissonmidpointmethodlangevin . This leads to improvement in the convergence rates compared to LMC under log concavity (see Table 1). he2021ergodicitybiasasymptoticnormality ; yu2023langevinmontecarlostrongly sharpen the bounds in shen2019randomizedmidpointmethodlogconcave .
kandasamy2024poissonmidpointmethodlangevin introduced the Poisson midpoint method for LMC (PLMC), a variant of RLMC, which converges whenever LMC converges, allowing analysis beyond log-concavity. PLMC gives a quadratic improvement in complexity in terms of $\epsilon$ when $\pi$ satisfies Logarithmic Sobolev Inequalities (LSI). In comparison, our work shows a cubic improvement for PLMC under log-concavity. Recently, altschuler2024shiftedcompositioniiilocal analyzed the randomized midpoint method for Overdamped LMC when the target satisfies LSI.
Note that the literature on MCMC considers 3 notions of convergence: KL-divergence, TV distance and 2-Wasserstein distance. In the case when $\pi$ is strongly log-concave, the Otto-Villani Theorem otto2000generalization shows that $\mathsf{KL}(\mu||\pi)\leq\epsilon^{2}\implies W_{2}^{2}(\mu,\pi)\lesssim\frac{\epsilon^{2}}{\alpha}$ and the Pinsker’s inequality shows that $\mathsf{KL}(\mu||\pi)\leq\epsilon^{2}\implies\mathsf{TV}(\mu,\pi)\lesssim\epsilon$. The condition of $\pi$ satisfying LSI is more general than log-concavity of the target. We refer to Table 1 for a detailed comparison of the results.
2.4 Our contributions
In the setting of sampling a log-concave target distribution $\pi(x)\propto\exp(-F(x))$:
(1) Overdamped PLMC achieves the desired error (1) after $\tilde{O}\bigr{[}\tfrac{\kappa^{4/3}+\kappa d^{1/3}}{\epsilon^{2/3}}\bigr{]}$ steps (Corollary 1).
(2) Underdamped PLMC achieves the same after $\tilde{O}\big{[}\frac{\kappa^{7/6}d^{1/6}}{\epsilon^{1/3}}+\frac{\kappa^{\frac{11p+6}{8p+6}}d^{\frac{p}{4p+3}}}{\epsilon^{\frac{p+2}{4p+3}}}\big{]}$ steps (Corollary 2). Here $p\in\mathbb{N}$ is arbitrary.
(3) To our knowledge, this is the state-of-the-art in $\epsilon$ dependence for discretization of Langevin dynamics, achieving a cubic improvement over LMC (Table 1).
Comparison to Strong Error Lower Bounds:
[…]
However, sampling algorithm guarantees consider ‘weak errors’ which are distances between $\mathsf{Law}(U_{T}(\omega))$ and $\mathsf{Law}(A(F,\omega,\tilde{\omega}))$. In particular, the Wasserstein-2 distance is the infimum of $L^{2}$ errors when $U_{T}$ is driven by $B_{t}(\omega)$ and $A(\cdot)$ queries $B_{t}^{\prime}(\omega)$ over all couplings of distinct Brownian motions $B_{t}(\omega)$ and $B^{\prime}_{t}(\omega)$. Our results for Poisson ULMC shows queries the oracle $\tilde{O}_{\kappa,d}(\frac{1}{\epsilon^{1/3}})$ times in expectation to achieve $W_{2}^{2}(\mathsf{Law}(A(F,\omega,\tilde{\omega})),\pi)\leq\frac{\epsilon^{2}d}{\alpha}$, a quadratic improvement.
We note that kandasamy2024poissonmidpointmethodlangevin obtained a complexity upper bound of $O_{d,\kappa}(\frac{1}{\sqrt{\epsilon}})$ for Underdamped ULMC under LSI assumptions for achieving $\mathsf{TV}\leq\epsilon$. The literature on sampling algorithms compares bounds of the form $W_{2}^{2}\leq\frac{\epsilon^{2}}{\alpha}$ to bounds of the form $\mathsf{TV}\leq\epsilon$ (see Section 2.3). Under this comparison our bound improves over prior art. However, we note that $\mathsf{TV}$ and $W_{2}^{2}$ bounds cannot be directly compared.
3 Results
We now present our main results.
Theorem 1.
Let $\tilde{X}_{tK+i}$ denote the iterates of Overdamped PLMC, and $X_{tK+i}$ the iterates of Overdamped LMC with stepsize $\eta/K.$ and $\mathsf{Law}(X_{0})=\mathsf{Law}(\tilde{X}_{0})$. Assume $\eta L\leq 1/8,$ and Assumption 1. Then there exist absolute constants $c_{1}$ and $c_{2}$ such that
$W_{2}^{2}(\mathsf{Law}(\tilde{X}_{tK}),\mathsf{Law}(X_{tK}))\lesssim$ $(\eta^{6}L^{4}dK+\eta^{4}L^{2}+\tfrac{\eta^{5}L^{4}}{\alpha})\cdot(Ldt+\tfrac{1}{\eta}\mathbb{E}(F(X_{0})-F(X_{tK}))$
$\tfrac{\eta^{3}L^{4}d}{\alpha^{2}}+\tfrac{\eta^{4}L^{4}d^{2}}{\alpha}+\exp(c_{1}d-(c_{2}\eta^{2}L^{2}K)^{-1})\cdot\tfrac{\eta^{2}L^{2}d}{\alpha}.$
The above theorem shows that $\tilde{X}_{tK}$ is close to $X_{tK}$ in Wasserstein-2 distance. However, running $tK$ iterations of PLMC requires only $O(t)$ gradient calls, as compared to $tK$ gradient calls for LMC. In the following corollary, we combine the Theorem 1 with the convergence results for $X_{tK+i}$ to $\pi$ given in durmus2018analysislangevinmontecarlo to deduce the convergence of $\tilde{X}_{tK+i}$. We refer to Section C.2 for its proof.
Corollary 1.
Let $\tilde{X}_{t}$ be the iterates of Overdamped PLMC as in Theorem 1. Assume the conditions from Theorem 1, and that $0<\epsilon\lesssim\min(1,\kappa^{-1/4},d^{-1/4}).$ Let $x^{*}$ be the unique minimizer of $F$, and let $\tilde{X}_{0}$ satisfy $\mathbb{E}[F(\tilde{X}_{0})-F(x^{*})]\leq C_{F}\kappa d$ for some $C_{F}>0$. Then there exist constants $C_{1},C_{2}>0$ depending only on $C_{F}$, $\log(\frac{W_{2}(U_{0},\pi)\sqrt{\alpha}}{\epsilon\sqrt{d}})$ and $\log(1/\epsilon),$ polynomially, such that if $\eta=C_{1}\min(\frac{\alpha^{1/3}\epsilon^{2/3}}{L^{4/3}},\frac{\epsilon^{2/3}}{d^{1/3}L}),$ $K=4\frac{\eta L}{\epsilon^{2}}$ and $N=C_{2}\big{[}\frac{\kappa^{4/3}+\kappa d^{1/3}}{\epsilon^{2/3}}\big{]}$. Then,
$W_{2}^{2}(\mathsf{Law}(\tilde{X}_{NK}),\pi)^{2}\leq\epsilon^{2}d/\alpha$
Remark 2.
The complexity bound for Overdamped LMC (durmus2018analysislangevinmontecarlo ) is $\tilde{O}(\kappa/\epsilon^{2})$ gradient calls, and that of Overdamped RLMC (yu2023langevinmontecarlostrongly ) is $\tilde{O}(\frac{\kappa}{\epsilon}+\frac{\kappa^{4/3}}{\epsilon^{2/3}})$ gradient calls. To our knowledge, our method is thus the best known discretization of overdamped Langevin dynamics, in terms of $\epsilon$ dependence. Note that assumption our on the initialization is very mild - $F$ can be optimized quickly under Assumption 1.
We now turn our attention to underdamped Langevin dynamics.
Theorem 3.
Let $\tilde{X}_{tK+i}$ denote the iterates of Underdamped PLMC, and $X_{tK+i}$ denote the iterates of Underdamped LMC with step-size $\eta/K$. Let $M=\begin{bmatrix}\mathbf{I}_{d}&0\\ \mathbf{I}_{d}&\frac{2}{\gamma}\mathbf{I}_{d}\end{bmatrix},$ and $p\geq 0$ be an integer. Assume $\eta/K\lesssim\frac{1}{\kappa\sqrt{L}},$ the conditions of Theorem 6, Assumption 1, and that $\gamma\eta<c_{0}$ for sufficiently small $c_{0}>0.$ Then with $\gamma=c\sqrt{L}$ for some $c\geq 2$, we have
$W_{2}^{2}(\mathsf{Law}(M\tilde{X}_{tK}),\mathsf{Law}(MX_{tK}))$ $\leq O\Big{[}\tfrac{\eta^{7}L^{9/2}d}{\alpha\gamma^{2}}t+\tfrac{\eta^{8}L^{4}d^{2}}{\gamma^{2}}t^{2}+\tfrac{\eta^{4p+4}K^{p-1}L^{2p+2}d^{p+1}}{\gamma^{2}}t^{p+1}\Big{]}$
$+\mathbb{E}[P(||V_{0}||,|F(\Psi_{0})-F(\Psi_{t})|^{+})],$
Where $P$ is a polynomial, and $\Psi$ is defined as $\Psi_{s}:=\tilde{U}_{sK}+\frac{1}{\gamma}\tilde{V}_{sK}.$ The complete bound is laid out in Section 6.1, for the sake of clarity.
Similar to Corollary 1, the following Corollary (proved in Section F) establishes complexity bounds.
Corollary 2.
Let ${\tilde{X}}_{tK+i}$ denote the iterates of Underdamped PLMC, as in Theorem 3. Assume the conditions of Theorem 3, and that $\epsilon\lesssim\min(1,\kappa d^{-1/2},\kappa^{\frac{p+3}{2(3p+1)}}d^{-\frac{p}{3p+1}})$. Set $K=\frac{\eta L\sqrt{2}}{0.94\epsilon\sqrt{\alpha}},$ $\gamma\asymp\sqrt{L}$, and initialize $V_{0}\sim\mathcal{N}(0,\mathbf{I}_{d}).$ Let $x^{*}$ be the unique minimizer of $F$, and initialize $X_{0}$ with $\mathbb{E}||X_{0}-x^{*}||^{2k}\lesssim d^{k}/L^{k}$ for $k=\max(2,p+1)$. Then there exist constants $C_{3},C_{4}>0$ depending only on $\log(\frac{W_{2}(U_{0},\pi)\sqrt{\alpha}}{\epsilon\sqrt{d}})$ and $\log(\frac{\kappa}{\epsilon}),$ polynomially, such that if $\eta=C_{3}\min\left(\frac{\epsilon^{1/3}}{\kappa^{1/6}d^{1/6}\sqrt{L}},\frac{\epsilon^{\frac{p+2}{4p+3}}}{\kappa^{\frac{3p}{8p+6}}d^{\frac{p}{4p+3}}\sqrt{L}}\right)$ and $N=C_{4}\big{[}\frac{\kappa^{7/6}d^{1/6}}{\epsilon^{1/3}}+\frac{\kappa^{\frac{11p+6}{8p+6}}d^{\frac{p}{4p+3}}}{\epsilon^{\frac{p+2}{4p+3}}}\big{]}$, we have
$W_{2}^{2}(\mathsf{Law}(\tilde{U}_{NK}),\pi)\leq\epsilon^{2}d/\alpha\,.$
Remark 4.
Our assumption on the initialization is very standard vempala2022rapidconvergenceunadjustedlangevin ; shen2019randomizedmidpointmethodlogconcave , and satisfied (for example) by $\mathcal{N}(x^{*},\mathbf{I}_{d}/L).$ With $p=3$, we get a complexity of $\tilde{O}(\frac{\kappa^{7/6}d^{1/6}+\kappa^{13/10}d^{1/5}}{\epsilon^{1/3}}).$ The complexity bound for Underdamped LMC is cheng2018-underdamped-lmc $\tilde{O}(\kappa^{3/2}/\epsilon)$, and that of Underdamped RLMC is $\tilde{O}(\frac{\kappa^{7/6}}{\epsilon^{1/3}}+\frac{\kappa}{\epsilon^{2/3}})$. To our knowledge, PLMC is the best known discretization of underdamped Langevin dynamics in terms of $\epsilon$.
4 Proof overview
Our proof follows from a coupling argument between the iterates of LMC with step-size $\eta/K$ (denoted by $X_{tK+i}$) and the iterates of PLMC with step-size $\eta$ (denoted by $\tilde{X}_{tK+i}$) to bound the $W_{2}$ distance. Considering the specific case of Overdamped LMC, we have:
$X_{tK+i+1}=T(X_{tK+i})+\sqrt{\tfrac{2\eta}{K}}Z_{tK+i};\quad\quad\tilde{X}_{tK+i+1}=T({\tilde{X}}_{tK+i})+\sqrt{\tfrac{2\eta}{K}}\tilde{Z}_{tK+i}.$ (4)
Here $T(x)=x-\frac{\eta}{K}\nabla F(x)$ denotes the gradient step, and $Z_{tK+i}$ denotes standard Gaussian noise. The term $\tilde{Z}_{tK+i}$ is a Gaussian perturbed with error induced by the Poisson midpoint method. Conditioned on $X_{tK+i}$ and ${\tilde{X}}_{tK+i}$, we provide a sharp bound on the $W_{2}$ distance between $Z_{tK+i}$ and $\tilde{Z}_{tK+i}$ using Lemma 1 (along with some additional calculations). The proof of this Lemma is in Section A.1. We note that this is similar to Lemma 7 of kandasamy2024poissonmidpointmethodlangevin , which was in turn adapted from zhai2017highdimensionalcltmathcalw2distance . The difference here is that our analysis avoids higher powers of $\nu,\beta$ in the bounds.
Lemma 1.
Let $V$ be a random vector in $\mathbb{R}^{d}$ satisfying the following conditions:
  1. 1.
    $||V||\leq\beta$ a.s., $\mathbb{E}[V]=0$, and $\mathbb{E}[VV^{T}]=\Sigma.$
  2. 2.
    $V$ lies in a one-dimensional subspace almost surely.
Let the random vector $Z\sim\mathcal{N}(0,\mathbf{I}_{d}),$ and independent of $V$. Let $\nu=\text{Tr}(\Sigma),$ Then,
$W_{2}^{2}(\mathsf{Law}(Z),\mathsf{Law}(Z+V))\leq\frac{11}{2}\nu^{2}+{\mathbf{1}}_{5\beta^{2}>1}\cdot 2\nu.$
Note that the bound $\nu^{2}$ can be much smaller than just the variance $\nu$ in our situation, leading to our sharp analysis. To describe our coupling (roughly), we take $X_{tK+i}$ and ${\tilde{X}}_{tK+i}$ to be $W_{2}$ optimally coupled. Conditioned on $X_{tK+i}$ and ${\tilde{X}}_{tK+i}$, we draw $Z_{tK+i}$ and $\tilde{Z}_{tK+i}$ to be conditionally $W_{2}$ optimally coupled with bounds on $W_{2}$ established in Lemma 1.
Under Assumption 1, the map $T$ is a contraction. Thus, applying this coupling to Equation (4), we conclude that (roughly, not exactly!) the following bound holds:
$W_{2}^{2}(X_{tK+i+1},\tilde{X}_{tK+i+1})\leq(1-\tfrac{\alpha\eta}{2K})W_{2}^{2}(\tilde{X}_{tK+i},X_{tK+i})+\tfrac{\eta^{4}L^{2}}{K}\mathbb{E}||\nabla F(\tilde{X}_{tK})||^{2}+\tfrac{\eta^{4}L^{4}d}{\alpha K}$
Unrolling the recursion above, we bound $W_{2}^{2}(\tilde{X}_{tK+i},X_{tK+i})$ in terms of $\sum_{t=0}^{N-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}$. We then show $\sum_{t=0}^{N-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}\lesssim LdN$ using the fact that LMC can be seen as gradient descent on $F$ with a very large Gaussian noise added. This allows us to prove Theorem 1. We then conclude Corollary 1 via standard convergence results of LMC to bound $W_{2}^{2}(X_{tK},\pi)$.
Underdamped PLMC
The proof in this case is similar to the overdamped case. Here we use the fact we can obtain a contraction map similar to $T$ in Equation (4) even for the underdamped case when we transform the co-ordinates $[u,v]^{T}$ under a suitable linear map (see (zhang2023improveddiscretizationanalysisunderdamped, , Lemma 16).
5 Proof for Overdamped PLMC
Let $X_{tK+i}$ denote the iterates of overdamped Langevin Monte-Carlo with step-size $\frac{\eta}{K}$. Let ${\tilde{X}}_{tK+i}$ denote the iterates of Poisson overdamped Langevin Monte-Carlo with step size $\eta$, and $\hat{\tilde{X}}_{tK+i}$ denote the cheap interpolants as defined in Section 2.
$X_{tK+i+1}$ $=X_{tK+i}-\tfrac{\eta}{K}\nabla F(X_{tK+i})+\sqrt{\tfrac{2\eta}{K}}Z_{tK+i}$
$\tilde{X}_{tK+i+1}$ $=\tilde{X}_{tK+i}+\eta H_{i,t}(\nabla F(\tilde{X}_{tk})-\nabla F(\hat{\tilde{X}}_{tK+i}))-\tfrac{\eta}{K}\nabla F(\tilde{X}_{tK})+\sqrt{\tfrac{2\eta}{K}}Y_{tK+i}$
$\hat{\tilde{X}}_{tK+i}$ $=\tilde{X}_{tK}-\tfrac{\eta i}{K}\nabla F(\tilde{X}_{tK})+\sqrt{\tfrac{2\eta}{K}}\sum_{j=0}^{i}Y_{tK+j}$
Here, the sequence $(Z_{tK+i})_{t,i}$ are i.i.d. $\mathcal{N}(0,\mathbf{I}_{d})$ and so are $(Y_{tK+i})_{t,i}$. ($X_{0}$, $(Z_{tK+i})_{t,i}$) and ($\tilde{X}_{0}$, $(H_{i,t})_{t,i}$$(Y_{tK+i})_{t,i}$) live on the same probability space with an arbitrary coupling. Define the event:
$\mathcal{G}=\{\tilde{X}_{tK}=y_{0},\tilde{X}_{tK+i}=y,\hat{\tilde{X}}_{tK+i}=\hat{y},X_{tK+i}=x\},$
with $x,y,\hat{y}$ and $y_{0}$ being arbitrary points in $\mathbb{R}^{d}.$ Let $\tilde{Z}_{tK+i}=\sqrt{\frac{\eta K}{2}}(H_{i,t}-1/K)(\nabla F(\tilde{X}_{tK})-\nabla F(\hat{\tilde{X}}_{tK+i}))+\sqrt{\frac{\eta}{2K}}(\nabla F(\tilde{X}_{tK+i})-\nabla F(\hat{\tilde{X}}_{tK+i}))+Y_{tK+i}.$ For any valid coupling of $X_{tK+i+1}$ and $\tilde{X}_{tK+i+1}$ conditioned on $\mathcal{G},$ the following holds.
Proposition 1.
Let Assumption 1 hold and let $\frac{\alpha\eta}{K}<1$. Let $V=\sqrt{\tfrac{\eta K}{2}}(H_{i,t}-1/K)[\nabla F(y_{0})-\nabla F(\hat{y})]$. Then we have,
$\mathbb{E}[||X_{tK+i+1}-\tilde{X}_{tK+i+1}||^{2}|\mathcal{G}]\leq(1-\tfrac{\alpha\eta}{2K})^{2}||x-y||^{2}+\tfrac{9\eta L^{2}}{\alpha K}||y-\hat{y}||^{2}+\tfrac{2\eta}{K}\mathbb{E}[||Y_{tK+i}+V-Z_{tK+i}||^{2}|\mathcal{G}].$
The proof of this Proposition is in section C.1. Let $\mathcal{G}^{\sigma}=\sigma(\tilde{X}_{tK},\tilde{X}_{tK+i},\hat{\tilde{X}}_{tK+i},X_{tK+i})$ be the sigma-algebra generated by the previous iterates. We apply Lemma 1 to bound the final term.
Corollary 3.
Let $\nu=\text{Tr}(VV^{T}|\mathcal{G})$ with $V$ defined as in Proposition 1, and $\beta^{2}=\frac{\eta KL^{2}}{2}||y_{0}-\hat{y}||^{2}.$ Let $S\in\mathcal{G}^{\sigma}$ be an event. Conditioned on $\mathcal{G}$, there exists a coupling of $(Y_{tK+i},H_{i,t})$ and $Z_{tK+i}$ such that
$\mathbb{E}[||Y_{tK+i}+V-Z_{tK+i}||^{2}|\mathcal{G}]\leq(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot 2\nu+\mathbf{1}_{S^{c}}\cdot\frac{11}{2}\nu^{2}.$
Proof.
Under the event $S$, we couple the Gaussians $Y_{tK+i}$ and $Z_{tK+i}$ identically (i.e, $Y_{tK+i}=Z_{tK+i}$). This gives $\mathbb{E}[\|Y_{tK+i}+V-Z_{tK+i}\|^{2}|\mathcal{G}]=\mathbb{E}[\|V\|^{2}|\mathcal{G}]=\nu.$ Under $S^{c},$ couple them as in the Lemma 1. ∎
Using the above results, we produce an explicit coupling of $X_{tK+i+1}$ and $\tilde{X}_{tK+i+1}$ to bound the Wasserstein distance between their distributions. We refer to Section C.3 for its proof.
Lemma 2.
Assume $\eta L/K\leq 1,$ and Assumption 1. Then there exist absolute constants $c_{1},c_{2}>0$ such that
$W_{2}^{2}(\mathsf{Law}(X_{tK+i+1}),\mathsf{Law}(\tilde{X}_{tK+i+1}))$ $\leq(1-\frac{\alpha\eta}{2K})W_{2}^{2}(\mathsf{Law}(\tilde{X}_{tK+i}),\mathsf{Law}(X_{tK+i}))+E,\text{ where}$
$E\lesssim(\eta^{6}L^{4}d+\tfrac{\eta^{4}L^{2}}{K}+\tfrac{\eta^{5}L^{4}}{\alpha K})\mathbb{E}||\nabla F(\tilde{X}_{tK})||^{2}+\tfrac{\eta^{4}L^{4}d}{\alpha K}+\tfrac{\eta^{5}L^{4}d^{2}}{K}+\exp(c_{1}d-(c_{2}\eta^{2}L^{2}K)^{-1})\cdot\tfrac{\eta^{3}L^{2}d}{K},$
5.1 Finishing the proof
We open the recursion in Lemma 2, summing the constant terms as a geometric series.
$W_{2}^{2}(X_{tK},\tilde{X}_{tK})$ $\lesssim\exp(-\alpha\eta t)W_{2}^{2}(X_{0},\tilde{X}_{0})^{2}+(\eta^{6}L^{4}Kd+{\eta^{4}L^{2}}+\tfrac{\eta^{5}L^{4}}{\alpha})\sum_{s=0}^{t-1}\mathbb{E}||\nabla F(\tilde{X}_{Ks})||^{2}$
$+\tfrac{\eta^{3}L^{4}d}{\alpha^{2}}+\tfrac{\eta^{4}L^{4}d^{2}}{\alpha}+\exp(c_{1}d-(c_{2}\eta^{2}L^{2}K)^{-1})\cdot\tfrac{\eta^{2}L^{2}d}{\alpha}.$
Note that $X_{0}=\tilde{X}_{0}$, so $W_{2}^{2}(X_{0},\tilde{X}_{0})=0.$ Moreover, $\sum_{t=0}^{N}\mathbb{E}||\nabla F(\tilde{X}_{tK})||^{2}$ is bounded in Lemma 9. Substituting this above completes the proof of Theorem 1.
6 Proof for Underdamped PLMC
Let $X_{tK+i}=[U_{tK+i},V_{tK+i}]^{T}$ denote the iterates of Underdamped LMC with stepsize $\eta/K$, and ${\tilde{X}}_{tK+i}=[{\tilde{U}}_{tK+i},{\tilde{V}}_{tK+i}]^{T}$ denote the iterates of Underdamped PLMC with step size $\eta$. We apply an affine transformation to the iterates, introduced in zhang2023improveddiscretizationanalysisunderdamped which makes Underdamped LMC a noisy contraction map (Lemma 16, zhang2023improveddiscretizationanalysisunderdamped ).
${\tilde{X}}_{tK+i}^{\prime}=M{\tilde{X}}_{tK+i},\text{ where }M=\begin{bmatrix}\mathbf{I}_{d}&0\\ \mathbf{I}_{d}&\frac{2}{\gamma}\mathbf{I}_{d}\end{bmatrix}.$
Denote $\tilde{W}_{tK+i}={\tilde{U}}_{tK+i}+\frac{2}{\gamma}{\tilde{V}}_{tK+i},$ so that ${\tilde{X}}_{tK+i}^{\prime}=[{\tilde{U}}_{tK+i},{\tilde{W}}_{tK+i}]^{T}.$ The transformed iterates satisfy the following recursion:
$\tilde{X}_{tK+i+1}^{\prime}=A_{\frac{\eta}{K}}^{\prime}{\tilde{X}}_{tK+i}^{\prime}+G_{\frac{\eta}{K}}^{\prime}b({\tilde{X}}_{tK+i}^{\prime})+\Gamma_{\frac{\eta}{K}}^{\prime}\tilde{Z}_{tK+i},$
where $b(X)=[-\nabla F(U),0]^{T},$ and
$A_{h}^{\prime}=\begin{bmatrix}\frac{1}{2}(1+\exp(-\gamma h))\mathbf{I}_{d}&\frac{1}{2}(1-\exp(-\gamma h))\mathbf{I}_{d}\\ \frac{1}{2}(1-\exp(-\gamma h))\mathbf{I}_{d}&\frac{1}{2}(1+\exp(-\gamma h))\mathbf{I}_{d}\end{bmatrix},\text{ }G_{h}^{\prime}=\begin{bmatrix}\frac{\gamma h-(1-\exp(-\gamma h))}{\gamma^{2}}\mathbf{I}_{d}&0\\ \frac{\gamma h+(1-\exp(-\gamma h))}{\gamma^{2}}\mathbf{I}_{d}&0\end{bmatrix},$
$\Gamma_{h}^{\prime 2}=\begin{bmatrix}\frac{4(1-\exp(-\gamma h)-(1-\exp(2\gamma h))+2\gamma h}{\gamma^{2}}\mathbf{I}_{d}&\frac{2\gamma h-(1-\exp(2\gamma h))}{\gamma^{2}}\mathbf{I}_{d}\\ \frac{2\gamma h-(1-\exp(2\gamma h))}{\gamma^{2}}\mathbf{I}_{d}&\frac{4(1-\exp(-\gamma h)+(1-\exp(2\gamma h))+2\gamma h}{\gamma^{2}}\mathbf{I}_{d}\end{bmatrix},$
And the noise is
$\tilde{Z}_{tK+i}=Y_{tK+i}+\Gamma_{\frac{\eta}{K}}^{\prime-1}G_{\frac{\eta}{K}}^{\prime}[b(\hat{\tilde{X}}_{tK+i}^{\prime})-b({\tilde{X}}_{tK+i}^{\prime})]+K(H_{i,t}-1/K)\Gamma_{\frac{\eta}{K}}^{\prime-1}G_{\frac{\eta}{K}}^{\prime}[b(\hat{\tilde{X}}_{tK+i}^{\prime})-b(\tilde{X}_{tK}^{\prime})],$
with $\hat{\tilde{X}}_{tK+i}^{\prime}=A_{\frac{i\eta}{K}}^{\prime}\tilde{X}_{tK}^{\prime}+G_{\frac{i\eta}{K}}^{\prime}b(\tilde{X}_{tK}^{\prime})+\sum_{j=0}^{i-1}A_{\frac{\eta(i-j-1)}{K}}^{\prime}\Gamma_{\tfrac{\eta}{K}}^{\prime}Y_{tK+j}.$ Our proof follows the same method as in the overdamped case. As before, $(Z_{tK+i})_{t,i}$ are i.i.d. $\mathcal{N}(0,\mathbf{I}_{d})$ and so are $(Y_{tK+i})_{t,i}$. ($X_{0}$, $(Z_{tK+i})_{t,i}$) and ($\tilde{X}_{0}$, $(H_{i,t})_{t,i}$$(Y_{tK+i})_{t,i}$) live on the same probability space with an arbitrary coupling. We apply will Lemma 5 to produce a particular coupling.
We condition on the previous iterates – with the following event:
$\mathcal{G}=\{\tilde{X}_{tK}^{\prime}=y_{0}=[u_{0},w_{0}]^{T},{\tilde{X}}_{tK+i}^{\prime}=y=[\tilde{u},\tilde{w}]^{T},\hat{\tilde{X}}_{tK+i}^{\prime}=\hat{y}=[\hat{u},\hat{w}]^{T},X_{tK+i}^{\prime}=x=[u,w]^{T}\},$
where $x,y,y_{0}$ and $\hat{y}$ are arbitrary points in $\mathbb{R}^{2d}$.
The following Proposition is proved in Section G.1. Writing the ULMC updates as $X^{\prime}_{tK+i+1}=T(X^{\prime}_{tK+i})+\Gamma^{\prime}_{\eta/K}Z_{tK+i}$, the proof of the following Proposition hinges on the fact that $T$ is a contraction mapping (i.e, Liphschitz constant $<1$).
Proposition 2.
Assume $\eta/K\lesssim\frac{1}{\kappa\sqrt{L}},$ $\frac{\gamma\eta}{K}<c_{0}$ for sufficiently small $c_{0}>0,$ and let $V=K(H_{i,t}-1/K)\Gamma_{\frac{\eta}{K}}^{\prime-1}G_{\frac{\eta}{K}}^{\prime}[b(\hat{\tilde{X}}_{tK+i})-b(\tilde{X}_{tK})].$ Then with $\gamma=c\sqrt{L}$ for some $c\geq 2,$ the following holds.
$\mathbb{E}[\|X_{tK+i+1}^{\prime}-\tilde{X}_{tK+i+1}^{\prime}\|^{2}|\mathcal{G}]$ $\leq(1-\Omega(\tfrac{\alpha\eta}{\gamma K}))\|x-y\|^{2}$
$+O\Big{[}\tfrac{\eta L^{2}}{\alpha\gamma K}\|\hat{u}-\tilde{u}\|^{2}+\tfrac{\eta}{\gamma K}\mathbb{E}[\|Y_{tK+i}+V-Z_{tK+i}\|^{2}|\mathcal{G}]\Big{]}.$
Having conditioned on $\mathcal{G}$, we can use 5 to bound $\mathbb{E}[\|Y_{tK+i}+V-Z_{tK+i}\|^{2}\mathcal{G}]$. We refer to Section E.1 for its proof.
Proposition 3.
Let $V$ be defined as in Proposition 2, and $p\geq 0$ be an integer. Conditioned on $\mathcal{G}$, there exists a coupling of $Y_{tK+i},H_{i,t}$ and $Z_{tK+i}$ such that
$\tfrac{\eta}{\gamma K}\mathbb{E}[\|Y_{tK+i}+V-Z_{tK+i}\|^{2}|\mathcal{G}]\lesssim\tfrac{\eta^{3}L^{4}}{\gamma^{3}K}\|u_{0}-\hat{u}\|^{4}+\tfrac{5^{p}\eta^{p+2}K^{p-1}L^{2p+2}}{\gamma^{p+2}}\|u_{0}-\hat{u}\|^{2p+2}.$
The proof of the following Lemma is given in Section E.2
Lemma 3.
Assume $\eta/K\lesssim\frac{1}{\kappa\sqrt{L}},$ and $\gamma\eta<c_{0}$ for sufficiently small $c_{0}>0.$ With $\gamma=c\sqrt{L}$ for some $c\geq 2$, the following is true.
$W_{2}^{2}$ $({X_{tK+i+1}^{\prime}},{\tilde{X}_{tK+i+1}^{\prime}})\leq(1-\Omega(\tfrac{\alpha\eta}{\gamma K}))W_{2}^{2}(X_{tK+i}^{\prime},{\tilde{X}}_{tK+i}^{\prime})+O\Big{[}\tfrac{\eta^{7}L^{4}}{\alpha\gamma K}\mathbb{E}\|\tilde{V}_{tK}\|^{2}$
$+\tfrac{\eta^{9}L^{4}}{\alpha\gamma K}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{2}+\tfrac{\eta^{8}L^{4}}{\alpha K}(d+\log K)+\tfrac{\eta^{7}L^{4}}{\gamma^{3}K}\mathbb{E}\|\tilde{V}_{tK}\|^{4}$
$+\tfrac{\eta^{11}L^{4}}{\gamma^{3}K}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{4}+\tfrac{\eta^{9}L^{4}}{\gamma K}(d+\log K)^{2}+\lambda_{p}[\tfrac{\eta^{3p+4}K^{p-1}L^{2p+2}}{\gamma^{p+2}}\mathbb{E}\|\tilde{V}_{tK}\|^{2p+2}$
$+\tfrac{\eta^{5p+6}K^{p-1}L^{2p+2}}{\gamma^{p+2}}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{2p+2}+\tfrac{\eta^{4p+5}K^{p-1}L^{2p+2}}{\gamma}(d+\log K)^{p+1}\Big{]},$
Where $\lambda_{p}$ is a constant depending only on $p.$
6.1 Finishing the proof
Open up the recursion, summing up the constant terms as a geometric series. This gives
$W_{2}^{2}({\mathsf{Law}(\tilde{X}_{tK}^{\prime})},$ ${\mathsf{Law}(X_{tK}^{\prime})})\lesssim\exp\Big{(}\Omega(-\tfrac{\alpha\eta t}{\sqrt{L}})\Big{)}W_{2}^{2}(\mathsf{Law}(X_{0}^{\prime}),\mathsf{Law}(\tilde{X}_{0}^{\prime}))$
$+\sum_{s=0}^{t}\Big{[}\tfrac{\eta^{7}L^{4}}{\alpha\gamma}\mathbb{E}\|\tilde{V}_{tK}\|^{2}+\tfrac{\eta^{9}L^{4}}{\alpha\gamma}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{2}\Big{]}+\tfrac{\eta^{7}L^{4}\gamma}{\alpha^{2}}(d+\log K)$
$+\sum_{s=0}^{t}\Big{[}\tfrac{\eta^{7}L^{4}}{\gamma^{3}}\mathbb{E}\|\tilde{V}_{tK}\|^{4}+\tfrac{\eta^{11}L^{4}}{\gamma^{3}}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{4}\Big{]}+\tfrac{\eta^{8}L^{4}}{\alpha}(d+\log K)^{2}$
$+\sum_{s=0}^{t}\lambda_{p}\Big{[}\tfrac{\eta^{3p+4}K^{p}L^{2p+2}}{\gamma^{p+2}}\mathbb{E}\|\tilde{V}_{tK}\|^{2p+2}+\tfrac{\eta^{5p+6}K^{p}L^{2p+2}}{\gamma^{p+2}}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{2p+2}\Big{]}$
$+\lambda_{p}\tfrac{\eta^{4p+4}K^{p}L^{2p+2}}{\alpha}(d+\log K)^{p+1}$
Note that $X_{0}^{\prime}=\tilde{X}_{0}^{\prime}$ by definition, so the first term is zero. The moments $\sum_{s=0}^{t}\mathbb{E}||\tilde{V}_{tK}||^{2p}$ and $\sum_{s=0}^{t}\mathbb{E}||\nabla F(\tilde{U}_{tK})||^{2p}$ are bounded in Theorem 6. Substitute them in, ignoring lower order terms via the assumption $\gamma\eta<c_{0}.$
$W_{2}^{2}($ ${\mathsf{Law}(\tilde{X}_{tK}^{\prime})},{\mathsf{Law}(X_{tK}^{\prime})})\lesssim\tfrac{\eta^{6}L^{4}}{\alpha\gamma^{2}}\left[\mathbb{E}\|\tilde{V}_{0}\|^{2}+\mathbb{E}|(F(\Psi_{0})-F(\Psi_{T}))^{+}+1\right]$
$+\tfrac{\eta^{7}L^{4}}{\alpha\gamma}t\left[\tfrac{\gamma^{2}}{L}+1\right](d+\log K)+\tfrac{\eta^{6}L^{4}}{\gamma^{4}}\left[\mathbb{E}\|\tilde{V}_{0}\|^{4}+\mathbb{E}|(F(\Psi_{0})-F(\Psi_{T}))^{+}|^{2}+1\right]$
$+\tfrac{\eta^{7}L^{4}}{\gamma^{3}}t\left[\tfrac{\gamma^{4}}{L^{2}}+(\gamma\eta t)\right](d+\log K)^{2}$
$+\lambda_{p}\tfrac{\eta^{3p+3}K^{p}L^{2p+2}}{\gamma^{p+3}}\left[\mathbb{E}\|\tilde{V}_{0}\|^{2p+2}+\mathbb{E}|(F(\Psi_{0})-F(\Psi_{T}))^{+}|^{p+1}+1\right]$
$+\lambda_{p}\tfrac{\eta^{3p+4}K^{p}L^{2p+2}}{\gamma^{p+2}}t\left[\tfrac{\gamma^{2p+2}}{L^{p+1}}+(\gamma\eta t)^{p}\right](d+\log K)^{p+1}$
$+\tfrac{\eta^{7}L^{4}\gamma}{\alpha^{2}}(d+\log K)+\tfrac{\eta^{8}L^{4}}{\alpha}(d+\log K)^{2}+\lambda_{p}\tfrac{\eta^{4p+4}K^{p}L^{2p+2}}{\alpha}(d+\log K)^{p+1}.$
7 Discussion and limitations:
Our work obtains novel convergence bounds for a randomized midpoint type method for the classical, well studied problem of strong log-concave sampling with Langevin dynamics. We improve the dependence on $\epsilon$ compared to all priors works and show that strong $L^{2}$ error lower bounds are not fundamental limits for sampling algorithms. However, the dependence on dimension $d$ and the condition number $\kappa$ in our bounds is slightly worse than prior works. We believe that further work is required to obtain better dependence on these parameters.
Acknowledgment:
We thank Sinho Chewi for pointing us to Cao_2021 , which presents lower bounds for the complexity of sampling via discretization of Underdamped Langevin Dynamics.
Appendix A General Technical Lemmas and Proofs
We will use Lemma 5, which is a specialization of Lemma 4. We check that the proof in [Lemma 1.6, [30]] does not require $n$ to be an integer and state the following:
Lemma 4 (Lemma 1.6, [30]).
Let $n>0$ and let $Y$ be an $\mathbb{R}^{k}$ valued random variable with mean 00, covariance $\Sigma/n$ and $\|Y\|\leq\frac{\beta}{\sqrt{n}}$ almost surely. For $t\geq 0$, let $Z_{t}$ denote a Gaussian of mean 0 and covariance $t\Sigma$ independent of $Y$. Let $\sigma_{\text{min}}^{2}$ denote the smallest eigenvalue of $\Sigma$. Then, for any $n\geq\frac{5\beta^{2}}{\sigma_{\text{min}}^{2}},$ we have
$W_{2}(Z_{1},Z_{1-1/n}+Y)\leq\frac{5\sqrt{k}\beta}{n\sqrt{n}}.$
We note that the following Lemma is similar in form and proof to Lemma 7 of [17].
Lemma 5.
Let $V$ be a random vector in $\mathbb{R}^{d}$ satisfying the following conditions:
  1. 1.
    $\|V\|\leq\beta$ a.s., $\mathbb{E}[V]=0$, and $\mathbb{E}[VV^{T}]=\Sigma.$
  2. 2.
    $V$ lies in a one-dimensional subspace almost surely.
Suppose the random vector $Z$ is distributed as $\mathcal{N}(0,\mathbf{I}_{d}),$ and independent of $V$. Let $\nu=\text{Tr}(\Sigma),$ Then,
$W_{2}^{2}((\sqrt{\mathbf{I}_{d}+\Sigma}Z),(Z+V))\leq 5\nu^{2}+{\mathbf{1}}_{5\beta^{2}>1}\cdot 2\nu.$
Proof.
These distributions are the same along all directions perpendicular to $V.$ We couple those directions identically. Let $V^{\prime}$ denote the projection of $V$ onto the direction spanned by itself, and $Z^{\prime}$ denote a one-dimensional Gaussian. We get
$W_{2}(\text{Law}(\sqrt{\mathbf{I}_{d}+\Sigma}Z),\text{Law}(Z+V))$ $\leq W_{2}({\text{Law}(\sqrt{1+\nu}Z^{\prime})},{\text{Law}(Z^{\prime}+V^{\prime})})$
$=\sqrt{1+\nu}W_{2}({\text{Law}(Z^{\prime})},{\text{Law}(\tfrac{Z^{\prime}}{\sqrt{1+\nu}}+\tfrac{V^{\prime}}{\sqrt{1+\nu}}})).$
Now set $k=1,$ $n=1+\frac{1}{\nu},$ and $\beta\to\beta\sqrt{n}$. Here $\sigma_{\text{min}}=1,$ which means $5\beta^{2}\leq 1$ is sufficient to apply Lemma 4.
$\mathbf{1}_{5\beta^{2}\leq 1}\cdot W_{2}^{2}(\text{Law}(\sqrt{\mathbf{I}_{d}+\Sigma}Z),\text{Law}(Z+V))$ $\leq\mathbf{1}_{5\beta^{2}\leq 1}\cdot\frac{25\beta^{2}\nu^{2}}{1+\nu}\leq 5\nu^{2}.$
When $5\beta^{2}>1,$ we couple $\mathsf{Law}(\sqrt{1+\nu}Z^{\prime})$ and $\mathsf{Law}(Z^{\prime}+V^{\prime})$ to have the same Gaussian noise $Z^{\prime}$, with $V^{\prime}$ sampled independently of $Z^{\prime}$. A simple computation yields
$\mathbf{1}_{5\beta^{2}>1}\cdot W_{2}^{2}(\text{Law}(\sqrt{1+\nu}Z^{\prime}),\text{Law}(Z^{\prime}+V^{\prime}))$ $\leq\mathbf{1}_{5\beta^{2}>1}\cdot 2\nu.$
A.1 Proof of Lemma 1
Let $Y$ and $Z$ be distributed as $\mathcal{N}(0,\mathbf{I}_{d}),$ and denote $Y^{\prime}=Y+V.$ Let $W$ be another random variable with distribution $\mathcal{N}(0,\mathbf{I}_{d}).$ Couple $Z$ and $Y$ to $\sqrt{1+\Sigma}W$ as follows:
  1. 1.
    Let $Q_{1}$ be the optimal $W_{2}$ coupling law between $\mathsf{Law}(Z)$ and $\mathsf{Law}(\sqrt{I+\Sigma}W)$ and $Q_{2}$ be the optimal $W_{2}$ coupling law between $\mathsf{Law}(Y^{\prime})$ and $\mathsf{Law}(\sqrt{I+\Sigma}W)$.
  2. 2.
    Sample $W$ from $\mathcal{N}(0,\mathbf{I}_{d}),$
  3. 3.
    Let $Q_{i}(\cdot|x)$ denote the law of the first co-ordinate under the joint distribution $Q_{i}$ when conditioned on the second co-ordinate being $x$. Conditioned on $W$, sample $(Z,Y^{\prime})$ from the product of the conditional laws of the optimal couplings: $Q_{1}(\cdot|\sqrt{I+\Sigma}W)\times Q_{2}(\cdot|\sqrt{I+\Sigma}W)$,
Notice that each of $Z$ and $Y^{\prime}$ are optimally coupled to $\sqrt{I+\Sigma}W$. Since $(a+b)^{2}\leq 2(a^{2}+b^{2}),$
$\mathbb{E}[\|Y+V-Z\|^{2}]$ $\leq 2\mathbb{E}[\|Y+V-\sqrt{\mathbf{I}_{d}+\Sigma}W\|^{2}]+2\mathbb{E}[\|\sqrt{\mathbf{I}_{d}+\Sigma}W-Z\|^{2}]$
$=2W_{2}^{2}(Y+V,\sqrt{\mathbf{I}_{d}+\Sigma}W)+2W_{2}^{2}(\sqrt{\mathbf{I}_{d}+\Sigma}W,Z).$
The former term is bounded in Lemma 5. The latter term is a Wasserstein distance between Gaussians, which has the following closed form.
$2W_{2}^{2}(\sqrt{\mathbf{I}_{d}+\Sigma}W,Z)^{2}=4+2\nu-4\sqrt{1+\nu}\leq\frac{1}{2}\nu^{2}.$
Appendix B Technical Results for OLMC
Lemma 6 (Lemma 12, [17]).
Let $M_{t,K}=\sup_{0\leq j<K}||\sum_{i=0}^{j}\sqrt{\frac{2\eta}{K}}Y_{tK+i}||$, and $p\in\mathbb{N}.$ Let $N_{t}:=\sum_{i=0}^{k-1}H_{i,t}.$ Then the following bounds are true.
$\sup_{0\leq i\leq k-1}||\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}||$ $\leq\eta||\nabla F(\tilde{X}_{tK})||+M_{t,k}.$
$\sup_{0\leq i\leq k-1}||\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK+i}||$ $\leq\eta LN_{t}\sup_{i\leq k-1}||\hat{\tilde{X}}_{tK+i}-\tilde{X}_{tK}||.$
$\mathbb{E}[M_{t,K}^{p}]$ $\leq(\eta d)^{p/2}.$
Lemma 7.
Let $\beta$ and $\nu$ be defined as in Lemma 3. Define the event $S\in\mathcal{G}^{\sigma}$ by $S=\{\frac{\eta^{4}L^{2}}{K}\|\nabla F(\tilde{X}_{tK})\|^{2}<\frac{\eta^{7}L^{4}}{K}\|\nabla F(\tilde{X}_{tK})\|^{4}\},$ we have
$\frac{\eta}{K}\mathbb{E}[(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot\nu+\mathbf{1}_{S^{c}}\cdot\nu^{2}]$ $\lesssim(\eta^{6}L^{4}d+\frac{\eta^{4}L^{2}}{K})\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}+\frac{\eta^{5}L^{4}d^{2}}{K}$
$+\exp(c_{1}d-(c_{2}\eta^{2}L^{2}K)^{-1})\cdot\frac{\eta^{3}L^{2}d}{K}$
where the expectation is taken over the distribution defined in the proof of 2.
Proof.
Since $H_{i,t}$ is a Bernoulli random variable with parameter $1/K,$ we have $\mathbb{E}[(H_{i,t}-1/K)^{2}]\leq 1/K.$ This gives us an upper bound on $\nu$, since $\nu=\mathbb{E}[\frac{\eta K}{2}(H_{i,t}-1/K)^{2}\|\nabla F(y_{0})-\nabla F(\hat{y})\|^{2}]\leq\frac{\eta L^{2}}{2}\|y_{0}-\hat{y}\|^{2}$ under assumption 1. This gives
$\frac{\eta}{K}(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot\nu+\mathbf{1}_{S^{c}}\cdot\nu^{2}\lesssim(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot\frac{\eta^{2}L^{2}}{K}\|y_{0}-\hat{y}\|^{2}+\mathbf{1}_{S^{c}}\cdot\frac{\eta^{3}L^{4}}{K}\|y_{0}-\hat{y}\|^{4}$
Now we apply item 1 of Lemma 6 to obtain the following.
$\frac{\eta^{2}L^{2}}{K}\|y_{0}-\hat{y}\|^{2}$ $\lesssim\frac{\eta^{4}L^{2}}{K}\|\nabla F(\tilde{X}_{tK})\|^{2}+\frac{\eta^{2}L^{2}}{K}M_{t,K}^{2}.$
$\frac{\eta^{3}L^{4}}{K}\|y_{0}-\hat{y}\|^{4}$ $\lesssim\frac{\eta^{7}L^{4}}{K}\|\nabla F(\tilde{X}_{tK})\|^{4}+\frac{\eta^{3}L^{4}}{K}M_{t,K}^{4}.$
Using $(\mathbf{1}_{5\beta^{2}>1}+\mathbf{1}_{S})\lesssim 1$ and $\mathbf{1}_{S^{c}}\frac{\eta^{7}L^{4}}{K}\|\nabla F(\tilde{X}_{tK})\|^{4}\leq\frac{\eta^{4}L^{2}}{K}\|\nabla F(\tilde{X}_{tK})\|^{2},$ we obtain
$\frac{\eta}{K}[(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot\nu+\mathbf{1}_{S^{c}}\cdot\nu^{2}]$ $\lesssim\frac{\eta^{4}L^{2}}{K}\|\nabla F(\tilde{X}_{tK})\|^{2}+\frac{\eta^{3}L^{4}}{K}M_{t,K}^{4}$
$+(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\frac{\eta^{2}L^{2}}{K}M_{t,K}^{2}.$
The mean second term is bounded in Lemma 6, and the expectation of the final term is bounded in Lemma 8. ∎
Lemma 8.
Let $\beta$ and $S$ be as in Lemma 7. There exists an absolute constants $c_{1}$ and $c_{2}$ such that
$\mathbb{E}[(\mathbf{1}_{5\beta^{2}>1}+\mathbf{1}_{S})\cdot\frac{\eta^{2}L^{2}}{K}M_{t,K}^{2}]\lesssim\eta^{6}L^{4}d\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}+\exp(c_{1}{d}-(c_{2}\eta^{2}L^{2}K)^{-1})\frac{\eta^{3}L^{2}d}{K}$
Proof.
Note that $S$ is independent of $M_{t,K},$ and by its definition we have $1_{S}\leq\eta^{3}L^{3}\|F(\tilde{X}_{tK})\|^{2}.$ As a result,
$\mathbb{E}[1_{S}\cdot\frac{\eta^{2}L^{2}}{K}M_{t,K}^{2}]\leq\eta^{3}L^{2}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}\cdot\mathbb{E}[\frac{\eta^{2}L^{2}}{K}M_{t,K}^{2}].$
Recall the definition of $\beta.$
$\beta$ $\leq\sqrt{\eta K}L\|\tilde{X}_{tK}-\hat{\tilde{X}}_{tK+i}\|$
$=\sqrt{\eta K}L\biggr{\|}\frac{\eta i}{K}\nabla F(\tilde{X}_{tK})+\sqrt{\frac{2\eta}{K}}\sum_{j=0}^{i}Y_{tK+j}\biggr{\|}.$
Applying triangle inequality and union bound, we get
$\mathbf{1}_{\sqrt{5}\beta>1}$ $\leq\mathbf{1}\{\sqrt{5}\eta^{3/2}K^{1/2}L\|\nabla F(\tilde{X}_{tK})\|>1\}+\mathbf{1}\{\sqrt{1}0\eta L\|\sum_{j=0}^{i}Y_{tK+j}\|>1\}.$
Note that $\tilde{X}_{tK}$ is independent of $M_{t,K}$. To handle the second term below, apply Cauchy Schwarz and a Gaussian concentration inequality.
$\mathbb{E}[\mathbf{1}_{5\beta^{2}>1}\cdot\frac{\eta^{2}L^{2}}{K}M_{t,K}^{2}]$ $\leq\mathbb{P}[\sqrt{5}\eta^{3/2}K^{1/2}L\|\nabla F(\tilde{X}_{tK})\|>1]\cdot\frac{\eta^{2}L^{2}}{K}\mathbb{E}[M_{t,K}^{2}]$
$+\mathbb{P}\big{[}\sqrt{1}0\eta L\|\sum_{j=0}^{i}Y_{tK+j}\|>1\big{]}^{1/2}\cdot\frac{\eta^{2}L^{2}}{K}\mathbb{E}[M_{t,K}^{4}]^{1/2}$
$\lesssim\eta^{3}KL^{2}\mathbb{E}[\|\nabla F(\tilde{X}_{tK})\|^{2}]\cdot\frac{\eta^{2}L^{2}}{K}\mathbb{E}[M_{t,K}^{2}]$
$+\exp(c_{1}d-(c_{2}\eta^{2}L^{2}K)^{-1})\cdot\frac{\eta^{2}L^{2}}{K}\mathbb{E}[M_{t,K}^{4}]^{1/2}.$
Where $c_{1},c_{2}>0$ are absolute constants. Applying item 2 of Lemma 6 completes the proof. ∎
Lemma 9.
Assuming $\eta L\leq 1/8,$ the following bound is true.
$\sum_{t=0}^{N-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}\lesssim\frac{1}{\eta}\mathbb{E}[F(\tilde{X}_{0})-F(\tilde{X}_{NK})]+{LdN}.$
Proof.
Since $F$ is smooth, we have (Lemma 3.4, [3])
$F(\tilde{X}_{(t+1)K})-F(\tilde{X}_{tK})\leq\langle\nabla F(\tilde{X}_{tK}),\tilde{X}_{(t+1)K}-\tilde{X}_{tK}\rangle+\frac{L}{2}\|\tilde{X}_{(t+1)K}-\tilde{X}_{tK}\|^{2}.$
By definition, $\tilde{X}_{(t+1)K}-\tilde{X}_{tK}=-\eta\nabla F(\tilde{X}_{tK})+\sum_{i=0}^{K-1}\eta H_{i,t}(\nabla F(\tilde{X}_{tK})-\nabla F(\hat{\tilde{X}}_{tK+i}))+\sum_{i=0}^{K-1}\sqrt{\frac{2\eta}{K}}Y_{tK+i}.$ Since $\mathbb{E}[H_{i,t}]=1/K$ and $\mathbb{E}[Y_{tK+i}]=0,$
$\mathbb{E}\langle\nabla F(\tilde{X}_{tK}),\tilde{X}_{(t+1)K}-\tilde{X}_{tK}\rangle$ $\leq-\eta\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}$
$+\sum_{i=0}^{K-1}\frac{\eta}{K}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|\cdot\|\nabla F(\tilde{X}_{tK})-\nabla F(\hat{\tilde{X}}_{tK+i})\|$
$\leq-\frac{\eta}{2}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}+\sum_{i=0}^{K-1}\frac{\eta}{2K}\mathbb{E}\|\nabla F(\tilde{X}_{tK})-\nabla F(\hat{\tilde{X}}_{tK+i})\|^{2}$
$\leq-\frac{\eta}{2}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}+\frac{\eta}{2}\sup_{0\leq i\leq K-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})-\nabla F(\hat{\tilde{X}}_{tK+i})\|^{2}$
$\leq-\frac{\eta}{2}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}+\eta^{3}L^{2}\|\nabla F(\tilde{X}_{tK})\|^{2}+\eta^{2}L^{2}d.$
Where in the second and final steps we used $ab\leq\frac{a^{2}+b^{2}}{2}$ and 6 respectively. Now we use $\|a+b\|^{2}\leq 2(\|a\|^{2}+\|b\|^{2})$ and $\mathbb{E}\|\sum_{i=0}^{K-1}\sqrt{\frac{2\eta}{K}}Y_{tK+i}\|^{2}=2\eta d$ to get
$\frac{L}{2}\|\tilde{X}_{(t+1)K}-\tilde{X}_{tK}\|^{2}\leq\eta^{2}L\|\nabla F(\tilde{X}_{tK})\|^{2}+\eta^{2}L\|\sum_{i=0}^{K-1}H_{i,t}(\nabla F(\tilde{X}_{tK})-\nabla F(\hat{\tilde{X}}_{tK+i}))\|^{2}+2\eta Ld.$
Let $N_{t}=\sum_{i=0}^{K-1}H_{i,t}$. Note that $\mathbb{E}[N_{t}^{2}]\leq 2,$ and $N_{t}$ is independent of $\tilde{X}_{tK}.$ Triangle inequality and 6 give
$\eta^{2}L\mathbb{E}\|\sum_{i=0}^{K-1}H_{i,t}(\nabla F(\tilde{X}_{tK})-\nabla F(\hat{\tilde{X}}_{tK+i}))\|^{2}$ $\leq\ \eta^{2}L\mathbb{E}[N_{t}\sup_{0\leq i\leq K-1}\mathbb{E}\|\nabla F(\tilde{X}_{tK})-\nabla F(\hat{\tilde{X}}_{tK+i})\|]^{2}$
$\leq 4\eta^{4}L^{3}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}+4\eta^{3}L^{3}d.$
Under our assumption $\eta L\leq 1/8,$ the terms $\eta^{3}L^{2}\|\nabla F(\tilde{X}_{tK})\|^{2},\eta^{4}L^{3}\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2},\eta^{2}L^{2}d$ and $\eta^{3}L^{3}d$ are negligible in order. Collecting the dominant terms, we get
$\eta\mathbb{E}\|\nabla F(\tilde{X}_{tK})\|^{2}\lesssim[F(\tilde{X}_{tK})-F(X_{(t+1)K})]+\eta Ld.$
This telescopes, leading to the result. ∎
Appendix C Deferred Proofs for OLMC
C.1 Proof of Proposition 1
Let $T(x)=x-\frac{\eta}{K}\nabla F(x).$ Under the assumption $\alpha\eta/K<1,$ it follows from the strong convexity and smoothness of $F$ that $T$ is a contraction with Lipschitz constant $(1-\frac{\alpha\eta}{K}).$ By definition, we have
$X_{tK+i+1}=T(X_{tK+i})+\sqrt{\frac{2\eta}{K}}Z_{tK+i},\text{ and }\tilde{X}_{tK+i+1}=T(\tilde{X}_{tK+i})+\sqrt{\frac{2\eta}{K}}\tilde{Z}_{tK+i}.$
Under the event $\mathcal{G}$, we have:
$\|X_{tK+i+1}-\tilde{X}_{tK+i+1}\|^{2}$ $=\|T(x)-T(y)\|^{2}+\frac{2\eta}{K}\|Z_{tK+i}-\tilde{Z}_{tK+i}\|^{2}$
$+2\sqrt{\frac{2\eta}{K}}\langle Z_{tK+i}-\tilde{Z}_{tK+i},T(x)-T(y)\rangle$
Conditioned on $\mathcal{G}$, $(H_{i,t}-1/K)$ has zero mean, and $Y_{tK+i}$, $Z_{tK+i}$ are standard Gaussians. This leads to
$\mathbb{E}[\|X_{tK+i+1}-\tilde{X}_{tK+i+1}\|^{2}|\mathcal{G}]$ $=\|T(x)-T(y)\|^{2}-{\frac{2\eta}{K}}\langle\nabla F(y)-\nabla F(\hat{y}),T(x)-T(y)\rangle$
$\quad+\frac{2\eta}{K}\mathbb{E}[\|Z_{tK+i}-\tilde{Z}_{tK+i}\|^{2}|\mathcal{G}]$
$\leq(1-\frac{\alpha\eta}{K})^{2}\|x-y\|^{2}+\frac{2\eta L}{K}(1-\frac{\alpha\eta}{K})\|y-\hat{y}\|\cdot\|x-y\|$
$\quad+\frac{2\eta}{K}\mathbb{E}[\|Z_{tK+i}-\tilde{Z}_{tK+i}\|^{2}|\mathcal{G}].$
The second term is bounded using the AM-GM inequality. For any arbitrary $\gamma>0,$
$\frac{2\eta L}{K}\|y-\hat{y}\|\cdot\|x-y\|\leq\frac{4\eta^{2}L^{2}}{\gamma}\|y-\hat{y}\|^{2}+\frac{\gamma}{K^{2}}\|x-y\|^{2}.$
In particular, with $\gamma=\alpha\eta K/2,$
$(1-\frac{\alpha\eta}{K})^{2}\|x-y\|^{2}+\frac{2\eta L}{K}(1-\frac{\alpha\eta}{K})\|y-\hat{y}\|\cdot\|x-y\|$
$\leq(1-\frac{\alpha\eta}{K})(1-\frac{\alpha\eta}{2K})\|x-y\|^{2}+(1-\frac{\alpha\eta}{K})\frac{8\eta L^{2}}{\alpha K}\|y-\hat{y}\|^{2}$
$\leq(1-\frac{\alpha\eta}{2K})^{2}\|x-y\|^{2}+\frac{8\eta L^{2}}{\alpha K}\|y-\hat{y}\|^{2}.$
By definition of $\tilde{Z}_{tK+i}$,
$\tilde{Z}_{tK+i}-Z_{tK+i}=\sqrt{\frac{\eta}{2K}}(\nabla F(y)-\nabla F(\hat{y}))+Y_{tK+i}+V-Z_{tK+i}.$
Square both sides, noting that $\mathbb{E}[Y_{tK+i}+V-Z_{tK+i}|\mathcal{G}]=0,$ and $\|\nabla F(y)-\nabla F(\hat{y})\|^{2}\leq L^{2}\|y-\hat{y}\|^{2}$ under assumption 1. This gives
$\frac{2\eta}{K}\mathbb{E}[\|Z_{tK+i}-\tilde{Z}_{tK+i}\|^{2}|\mathcal{G}]$ $=\frac{\eta^{2}L^{2}}{K^{2}}\|y-\hat{y}\|^{2}+\frac{2\eta}{K}\mathbb{E}[\|Y_{tK+i}+v-Z_{tK+i}\|^{2}|\mathcal{G}]$
$\leq\frac{\eta L^{2}}{\alpha K}\|y-\hat{y}\|^{2}+\frac{2\eta}{K}\mathbb{E}[\|Y_{tK+i}+v-Z_{tK+i}\|^{2}|\mathcal{G}].$
C.2 Proof of Corollary 1
Proof.
Wasserstein-2 is a metric on the space of probability measures over $\mathbb{R}^{d}$ with finite second moments. By triangle inequality
$W_{2}^{2}(\mathsf{Law}(\tilde{X}_{NK}),\pi)$ $\leq(W_{2}(\mathsf{Law}(\tilde{X}_{NK}),\mathsf{Law}(X_{NK}))+W_{2}(\mathsf{Law}(X_{NK}),\pi))^{2}$
$\lesssim W_{2}^{2}(\mathsf{Law}(\tilde{X}_{NK}),\mathsf{Law}(X_{NK}))+W_{2}^{2}(\mathsf{Law}(X_{NK}),\pi).$
We show under the conditions of our Corollary that each of these terms is $O(\epsilon^{2}d/\alpha).$ To deal with the second term, recall the following Theorem for the convergence of Langevin Monte-Carlo.
Theorem 5 (Corollary 10, [9]).
Suppose Assumption 1 is true. Let $X_{n}$ denote the iterates of Langevin Monte-Carlo with step-size $\gamma_{\epsilon}.$ Then, with
$\gamma_{\epsilon}=\frac{\epsilon^{2}}{4L},\;\;\;\;n_{\epsilon}\geq\lceil\log(\frac{2W_{2}^{2}(X_{0},\pi)\alpha}{\epsilon^{2}d})\gamma_{\epsilon}^{-1}\alpha^{-1}\rceil$
we have $W_{2}^{2}(X_{n_{\epsilon}},\pi)\leq\frac{\epsilon^{2}d}{\alpha}.$
By our choice of $K$, we have $\frac{\eta}{K}=\frac{\epsilon^{2}}{4L},$ and by the above Theorem, we have $W_{2}^{2}(X_{NK},\pi)\leq\frac{\epsilon^{2}d}{\alpha}$ for $N=\log(\frac{2W_{2}^{2}(X_{0},\pi)\alpha}{\epsilon^{2}d})(\alpha\eta)^{-1}.$ Let $L_{1}=2\max(C_{F},\log(\frac{2W_{2}^{2}(X_{0},\pi)\alpha}{\epsilon^{2}d})).$ Now apply Theorem 1 with
$\eta\asymp\min\Big{(}\frac{\epsilon^{2/3}}{L_{1}^{1/3}L},\frac{\epsilon^{1/2}}{\kappa^{1/4}L_{1}^{1/4}L},\frac{\epsilon^{2/3}}{d^{1/6}L_{1}^{1/6}L},\frac{\epsilon^{2/3}}{\kappa^{1/3}L},\frac{\epsilon^{1/2}}{d^{1/4}L},(\frac{c_{2}\epsilon^{2}}{c_{1}d-\log\epsilon^{2}})^{1/3}\cdot\frac{1}{L}\Big{)}$
and $N$ as above, to see $W_{2}^{2}(\mathsf{Law}(\tilde{X}_{NK}),\mathsf{Law}(X_{NK}))\lesssim\frac{\epsilon^{2}d}{\alpha}.$ We assume
$\epsilon\lesssim\min\Big{(}1,\kappa^{-1/6}L_{1}^{-1/6},d^{-1/8}L_{1}^{-1/8},\kappa^{-1/4},d^{-1/4}\Big{)},$
which is a regime in which the batch size $K=\frac{4\eta L}{\epsilon^{2}}\geq 1.$
C.3 Proof of Lemma 2
Proof.
Recall that given $y_{0},y,\hat{y},x\in\mathbb{R}^{d}$: $\mathcal{G}:=\{\tilde{X}_{tK}=y_{0},\tilde{X}_{tK+i}=y,\hat{\tilde{X}}_{tK+i}=\hat{y},X_{tK+i}=x\}$. By definition, conditioned on $\mathcal{G}$, we have:
$X_{tK+i+1}=x-\frac{\eta}{K}\nabla F(x)+\sqrt{\frac{2\eta}{K}}Z_{tK+i}$
$\tilde{X}_{tK+i+1}=y+\eta H_{i,t}(\nabla F(y_{0})-\nabla F(\hat{y}))-\frac{\eta}{K}\nabla F(y_{0})+\sqrt{\frac{2\eta}{K}}Y_{tK+i}.$
Conditioned on $\mathcal{G}$, we couple $(Y_{tK+i},H_{i,t})$ and $Z_{tK+i}$ by choosing the coupling produced in Corollary 3. This allows us to define $(X_{tK+i+1},\tilde{X}_{tK+i+1})$ using the equations above, conditioned on $\mathcal{G}$. This gives a conditional coupling of $(Y_{tK+i},H_{i,t},Z_{tK+i},X_{tK+i+1},\tilde{X}_{tK+i+1})$.
We produce an unconditional coupling as follows: Couple $X_{tK+i}$ and ${\tilde{X}}_{tK+i}$ optimally w.r.t. to $W_{2}$, then sample $\hat{\tilde{X}}_{tK+i}$ and $\tilde{X}_{tK}$ jointly conditioned on ${\tilde{X}}_{tK+i}$. Conditioned on $(\hat{\tilde{X}}_{tK+i},\tilde{X}_{tK},X_{tK+i},{\tilde{X}}_{tK+i})$ (i.e, $\mathcal{G}^{\sigma}$), we then sample $(Z_{tK+i},Y_{tK+i},H_{i,t},X_{tK+i+1},\tilde{X}_{tK+i+1})$ from the conditional coupling described above. Taking the expectation in Proposition 1, and using the bounds in Corollary 3 we get:
$W_{2}^{2}(X_{tK+i+1},\tilde{X}_{tK+i+1})^{2}\leq(1-\frac{\alpha\eta}{2K})^{2}W_{2}^{2}(X_{tK+i},{\tilde{X}}_{tK+i})^{2}+E,$
where $E\lesssim\frac{\eta L^{2}}{\alpha K}\mathbb{E}||{\tilde{X}}_{tK+i}-\hat{\tilde{X}}_{tK+i}||^{2}+\frac{\eta}{K}\mathbb{E}[(\mathbf{1}_{S}+\mathbf{1}_{5\beta^{2}>1})\cdot 2\nu+\mathbf{1}_{S^{c}}\cdot\frac{11}{2}\nu^{2}]$ and $S\in\mathcal{G}^{\sigma}$ is any event. We choose a particular event $S$ and bound the latter term in Lemma 7. The former term is bounded below, using items 1 and 2 of Lemma 6.
$\frac{\eta L^{2}}{\alpha K}\mathbb{E}||{\tilde{X}}_{tK+i}-\hat{\tilde{X}}_{tK+i}||^{2}\lesssim\frac{\eta^{3}L^{4}}{\alpha K}\mathbb{E}\left[N_{t}^{2}\sup_{j\leq K-1}||\tilde{X}_{tK}-\hat{\tilde{X}}_{tK+j}||^{2}\right]$
Note that $N_{t}$ is independent of $y_{0}$ and $\hat{y},$ and $\mathbb{E}[N_{t}^{2}]\lesssim 1.$ Along with item 2 of Lemma 6, this gives
$\frac{\eta L^{2}}{\alpha K}\mathbb{E}||{\tilde{X}}_{tK+i}-\hat{\tilde{X}}_{tK+i}||^{2}$ $\lesssim\frac{\eta^{5}L^{4}}{\alpha K}\mathbb{E}||\nabla F(\tilde{X}_{tK})||^{2}+\frac{\eta^{3}L^{4}}{\alpha K}\mathbb{E}[M_{t,K}^{2}]$
$\lesssim\frac{\eta^{5}L^{4}}{\alpha K}\mathbb{E}||\nabla F(\tilde{X}_{tK})||^{2}+\frac{\eta^{4}L^{4}d}{\alpha K}.$
Appendix D Technical Results for ULMC
Lemma 10 (Lemma 21, [17]).
Let $\pi$ denote projection onto the position axis: $\pi[u,v]^{T}=[u,0]^{T}.$ Let $M_{t,K}=\sup_{0\leq i<K}\|\sum_{j=0}^{i}A_{\frac{\eta(i-j)}{K}}\Gamma_{\frac{\eta}{K}}Z_{tK+j}\|_{\pi}.$ Then the following inequalities are true.
$\|\hat{\tilde{U}}_{tK+i}-\tilde{U}_{tK}\|$ $\lesssim\eta\|\tilde{V}_{tK}\|+\eta^{2}\|\nabla F(\tilde{U}_{tK})\|+M_{t,K}$
$\mathbb{E}[M_{t,K}^{p}]$ $\lesssim\exp(\frac{p\gamma\eta}{2})\gamma^{p/2}\eta^{3p/2}(d+\log K)^{p/2}.$
Lemma 11 (Lemmas 18/19, [17]).
[…]
  1. 1.
    $\|\Pi A_{\frac{j\eta}{K}}G_{\frac{\eta}{K}}\|\lesssim\frac{\eta^{2}}{K}$
  2. 2.
    $\|\hat{\tilde{X}}_{tK+i}-{\tilde{X}}_{tK+i}\|_{\mathbb{R}^{2d}}\leq\sum_{j=0}^{i-1}KH_{j,t}\|A_{\frac{(i-j-1)\eta}{K}}G_{\frac{\eta}{K}}(b(\hat{\tilde{X}}_{tK+i})-b(\tilde{X}_{tK}))\|_{\mathbb{R}^{2d}}.$
Theorem 6 (Theorem 4, [17]).
Fix $p\geq 1,$ and let $\mathcal{S}_{2p}(\nabla F)=\sum_{t=0}^{T}\|\nabla F(\tilde{U}_{tK})\|^{2p}$. Let $\mathcal{S}_{2p}(V)=\sum_{t=0}^{T}\|\tilde{V}_{tK}\|^{2p},$ and $\Psi_{t}=\tilde{U}_{tK}+\frac{1}{\gamma}\tilde{V}_{tK}.$ There exist constants $C_{p},c_{p},\bar{c}_{p}>0$ such that whenever: $\gamma\geq C_{p}\sqrt{L}$, $\alpha\gamma<c_{p}$, $\frac{\eta^{3p-1}T^{p-1}L^{2p}}{\gamma^{p+1}}<\bar{c}_{p}$, the following results hold:
$\mathcal{S}_{2p}(\nabla F)$ $\leq C_{p}\frac{\gamma^{2p-1}}{\eta}\left[\mathbb{E}\|\tilde{V}_{0}\|^{2p}+\mathbb{E}|(F(\Psi_{0})-F(\Psi_{T}))^{+}|^{p}+1\right]+$
$\quad C_{p}T\left[\tfrac{\gamma^{4p}}{L^{p}}+(\gamma\eta T)^{p-1}\gamma^{2p}\right](d+\log K)^{p}$
$\mathcal{S}_{2p}(V)$ $\leq C_{p}\frac{1}{\gamma\eta}\left[\mathbb{E}\|\tilde{V}_{0}\|^{2p}+\mathbb{E}|(F(\Psi_{0})-F(\Psi_{T}))^{+}|^{p}+1\right]$
$\quad+C_{p}T\left[\tfrac{\gamma^{2p}}{L^{p}}+(\gamma\eta T)^{p-1}\right](d+\log K)^{p}$
Appendix E Deferred Proofs for ULMC
E.1 Proof of Proposition 3
Proof.
Recall that $V=K(H_{i,t}-1/K)\Gamma_{\frac{\eta}{K}}^{\prime-1}G_{\frac{\eta}{K}}^{\prime}[b(\hat{\tilde{X}}_{tK+i}^{\prime})-b(\tilde{X}_{tK}^{\prime})],$ and let $\beta=\sqrt{\frac{\eta K}{\gamma}}L\|u_{0}-\hat{u}\|.$ We know $H_{i,t}\leq 1$ since it is a Bernoulli random variable, and $\|b(\hat{\tilde{X}}_{tK+i}^{\prime})-b(\tilde{X}_{tK}^{\prime})\|\leq L\|\hat{u}-u_{0}\|$ under Assumption 1. By Proposition 4, we have $\|G_{\frac{\eta}{K}}^{\prime T}\Gamma_{\frac{\eta}{K}}^{\prime-2}G_{\frac{\eta}{K}}^{\prime}\|\leq\frac{\eta}{\gamma K}.$ It follows that $\|V\|^{2}\leq\beta^{2}$. Now let $\nu=\text{Tr}(\mathbb{E}[VV^{T}|\mathcal{G}]).$ Since $\mathbb{E}[(H_{i,t}-1/K)^{2}]\leq 1/K,$ it follows that $\nu\leq\frac{\eta L^{2}}{\gamma}\|u_{0}-\hat{u}\|^{2}.$ Applying Lemma 1 gives
$\mathbb{E}[\|Y_{tK+i}+V-Z_{tK+i}\|^{2}|\mathcal{G}]$ $\lesssim\frac{\eta^{2}L^{4}}{\gamma^{2}}\|u_{0}-\hat{u}\|^{4}+\mathbf{1}_{5\beta^{2}>1}\cdot\frac{\eta L^{2}}{\gamma}\|u_{0}-\hat{u}\|^{2}$
$\lesssim\frac{\eta^{2}L^{4}}{\gamma^{2}}\mathbb{E}\|u_{0}-\hat{u}\|^{4}+\frac{5^{p}\eta^{p+1}K^{p}L^{2p+2}}{\gamma^{p+1}}\mathbb{E}\|u_{0}-\hat{u}\|^{2p+2}.$
In the last line, we have used $\mathbf{1}_{5\beta^{2}>1}\leq(5\beta^{2})^{p}=\frac{5^{p}\eta^{p}K^{p}L^{2p}}{\gamma^{p}}\|u_{0}-\hat{u}\|^{2p}.$ Multiplying this inequality by $\frac{\eta}{\gamma K}$ finishes the proof. ∎
E.2 Proof of Lemma 3
Proof.
Recall that given $y_{0},y,\hat{y},x\in\mathbb{R}^{d}$:
$\mathcal{G}:=\{\tilde{X}_{tK}^{\prime}=y_{0}=[u_{0},v_{0}]^{T},{\tilde{X}}_{tK+i}^{\prime}=y=[\tilde{u},\tilde{v}]^{T},\hat{\tilde{X}}_{tK+i}^{\prime}=\hat{y}=[\hat{u},\hat{v}]^{T},X_{tK+i}^{\prime}=x=[u,v]^{T}\}.$
By definition, conditioned on $\mathcal{G}$, we have
$X_{tK+i+1}^{\prime}=A_{\frac{\eta}{K}}^{\prime}x+G_{\frac{\eta}{K}}^{\prime}b(x)+\Gamma_{\frac{\eta}{K}}^{\prime}Z_{tK+i},$
$\tilde{X}_{tK+i+1}^{\prime}$ $=A_{\frac{\eta}{K}}^{\prime}y+G_{\frac{\eta}{K}}^{\prime}b(y)+\Gamma_{\frac{\eta}{K}}^{\prime-1}G_{\frac{\eta}{K}}^{\prime}[b(\hat{y})-b(y)]$
$+K(H_{i,t}-1/K)\Gamma_{\frac{\eta}{K}}^{\prime-1}G_{\frac{\eta}{K}}^{\prime}[b(\hat{y})-b(y_{0})]+\Gamma_{\frac{\eta}{K}}^{\prime}Y_{tK+i}.$
Conditioned on $\mathcal{G}$ we couple $Y_{tK+i},H_{i,t}$ and $Z_{tK+i}$ by choosing the coupling produced in Lemma 3. This allows us to define the distribution of $(X_{tK+i+1}^{\prime},\tilde{X}_{tK+i+1}^{\prime})$ using the equations above, conditioned on $\mathcal{G}$, and gives us a conditional coupling of $(Y_{tK+i},H_{i,t},Z_{tK+i},X_{tK+i+1}^{\prime},\tilde{X}_{tK+i+1}^{\prime})$ given $\mathcal{G}$.
We produce an unconditional coupling as follows. Couple $X_{tK+i}^{\prime}$ and ${\tilde{X}}_{tK+i}^{\prime}$ optimally w.r.t. $W_{2}$, then sample $\hat{\tilde{X}}_{tK+i}^{\prime}$ and $\tilde{X}_{tK}^{\prime}$ jointly conditioned on ${\tilde{X}}_{tK+i}^{\prime}.$ Conditioned on $(X_{tK+i}^{\prime},{\tilde{X}}_{tK+i}^{\prime},\hat{\tilde{X}}_{tK+i}^{\prime},\tilde{X}_{tK}^{\prime})$ we then sample $(Z_{tK+i},Y_{tK+i},H_{i,t},X_{tK+i+1}^{\prime},\tilde{X}_{tK+i+1})$ using the conditional coupling described above. We now take the expectation in Proposition 2, after substituting the bound in Proposition 3. This gives
$W_{2}^{2}({X_{tK+i+1}^{\prime}},$ ${\tilde{X}_{tK+i+1}^{\prime}})\leq(1-\Omega(\frac{\alpha\eta}{\gamma K}))W_{2}^{2}(X_{tK+i}^{\prime},{\tilde{X}}_{tK+i}^{\prime})+E,\text{ where}$
$E$ $\lesssim\frac{\eta L^{2}}{\alpha\gamma K}\mathbb{E}\|\hat{\tilde{U}}_{tK+i}-{\tilde{U}}_{tK+i}\|^{2}+\frac{\eta^{3}L^{4}}{\gamma^{3}K}\mathbb{E}\|\hat{\tilde{U}}_{tK+i}-\tilde{U}_{tK}\|^{4}$
$+\frac{5^{p}\eta^{p+2}K^{p-1}L^{2p+2}}{\gamma^{p+2}}\mathbb{E}\|\hat{\tilde{U}}_{tK+i}-\tilde{U}_{tK}\|^{2p+2}.$
We now bound each of the error terms individually. Recall $N_{t}:=\sum_{i=0}^{K-1}H_{i,t}$ and let $M_{t,K}$ be as defined in Lemma 10.
$\frac{\eta L^{2}}{\alpha\gamma K}\mathbb{E}\|\hat{\tilde{U}}_{tK+i}-{\tilde{U}}_{tK+i}\|^{2}$ $\leq\frac{\eta L^{2}}{\alpha\gamma K}\mathbb{E}\Big{[}\sum_{j=0}^{i-1}KH_{j,t}\|A_{\frac{(i-1-j)\eta}{K}}G_{\frac{\eta}{K}}(b(\hat{\tilde{X}}_{tK+j})-b(\tilde{X}_{tK}))\|\Big{]}^{2}$
$\lesssim\frac{\eta L^{2}}{\alpha\gamma K}\mathbb{E}\Big{[}\sum_{j=0}^{i-1}H_{j,t}\cdot{\eta^{2}L}\|\hat{\tilde{U}}_{tK+j}-\tilde{U}_{tK}\|\Big{]}^{2}$
$\lesssim\frac{\eta^{5}L^{4}}{\alpha\gamma K}\mathbb{E}\Big{[}N_{t}^{2}\sup_{0\leq j<K}\|\hat{\tilde{U}}_{tK+j}-\tilde{U}_{tK}\|^{2}\Big{]}$
$\lesssim\frac{\eta^{5}L^{4}}{\alpha\gamma K}\mathbb{E}\Big{[}\sup_{0\leq j<K}\|\hat{\tilde{U}}_{tK+j}-\tilde{U}_{tK}\|^{2}\Big{]}$
$\lesssim\frac{\eta^{7}L^{4}}{\alpha\gamma K}\mathbb{E}\|\tilde{V}_{tK}\|^{2}+\frac{\eta^{9}L^{4}}{\alpha\gamma K}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{2}+\frac{\eta^{5}L^{4}}{\alpha\gamma K}\mathbb{E}[M_{t,K}^{2}]$
$\lesssim\frac{\eta^{7}L^{4}}{\alpha\gamma K}\mathbb{E}\|\tilde{V}_{tK}\|^{2}+\frac{\eta^{9}L^{4}}{\alpha\gamma K}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{2}+\frac{\eta^{8}L^{4}}{\alpha K}(d+\log K).$
In the first inequality, we have used item 2 of Lemma 11. In the second, we have used item 1 of Lemma 11 and Assumption 1. In the fourth we have used that $N_{t}$ is independent of the iterates, and $\mathbb{E}[N_{t}]^{2}\lesssim 1.$ In the fifth and last inequalities, we have used items 1 and 2 of Lemma 10 respectively, with the assumption that $\gamma\eta$ is bounded.
$\frac{\eta^{3}L^{4}}{\gamma^{3}K}\mathbb{E}\|\hat{\tilde{U}}_{tK+i}-$ $\tilde{U}_{tK}\|^{4}\lesssim\frac{\eta^{7}L^{4}}{\gamma^{3}K}\mathbb{E}\|\tilde{V}_{tK}\|^{4}+\frac{\eta^{11}L^{4}}{\gamma^{3}K}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{4}+\frac{\eta^{3}L^{4}}{\gamma^{3}K}\mathbb{E}[M_{t,K}^{4}]$
$\lesssim\frac{\eta^{7}L^{4}}{\gamma^{3}K}\mathbb{E}\|\tilde{V}_{tK}\|^{4}+\frac{\eta^{11}L^{4}}{\gamma^{3}K}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{4}+\frac{\eta^{9}L^{4}}{\gamma K}(d+\log K)^{2}.$
The above inequality follows from items 1 and 2 of Lemma 10, with the assumption that $\gamma\eta$ is bounded. Now, for some constant $\lambda_{p}$ depending only on $p$:
$\frac{5^{p}\eta^{p+2}K^{p-1}L^{2p+2}}{\gamma^{p+2}}$ $\mathbb{E}\|\hat{\tilde{U}}_{tK+i}-\tilde{U}_{tK}\|^{2p+2}\leq\lambda_{p}^{\prime}\Big{[}\frac{\eta^{3p+4}K^{p-1}L^{2p+2}}{\gamma^{p+2}}\mathbb{E}\|\tilde{V}_{tK}\|^{2p+2}$
$+\frac{\eta^{5p+6}K^{p-1}L^{2p+2}}{\gamma^{p+2}}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{2p+2}+\frac{\eta^{p+2}K^{p-1}L^{2p+2}}{\gamma^{p+2}}\mathbb{E}[M_{t,K}^{2p+2}]\Big{]}$
$\leq\lambda_{p}\Big{[}\frac{\eta^{3p+4}K^{p-1}L^{2p+2}}{\gamma^{p+2}}\mathbb{E}\|\tilde{V}_{tK}\|^{2p+2}+\frac{\eta^{5p+6}K^{p-1}L^{2p+2}}{\gamma^{p+2}}\mathbb{E}\|\nabla F(\tilde{U}_{tK})\|^{2p+2}$
$+\frac{\eta^{4p+5}K^{p-1}L^{2p+2}}{\gamma}(d+\log K)^{p+1}\Big{]}.$
As before, the above inequality follows from items 1 and 2 of Lemma 10. ∎
Appendix F Proof of Corollary 2
Proof.
The assumption $\epsilon\lesssim\min(1,\kappa d^{-1/2},\kappa^{\frac{p+3}{2(3p+1)}}d^{-\frac{p}{3p+1}})$ ensures that the batch size $K\asymp\frac{\eta L}{\epsilon\sqrt{\alpha}}\geq 1$ making our algorithm valid. Since $W_{2}$ is a metric, triangle inequality gives
$W_{2}^{2}(\tilde{U}_{NK},\pi)\lesssim W_{2}^{2}(\tilde{U}_{NK},U_{NK})+W_{2}^{2}(U_{NK},\pi)\leq W_{2}^{2}(\tilde{X}_{NK},X_{NK})+W_{2}^{2}(U_{NK},\pi).$
Under the conditions of the Corollary, we show that both these terms are $\lesssim\frac{\epsilon^{2}d}{\alpha}.$ Recall the following Theorem for the convergence of Underdamped LMC.
Theorem 7 (Corollary of Theorem 2, [7]).
Let $F$ satisfy Assumption 1. In addition, let the initial condition of ULMC be drawn from the product distribution $\mu=\mathcal{N}(0,\mathbf{I}_{d})\otimes\nu_{0}.$ For $\gamma=c\sqrt{L}$ and step-size $h=\frac{0.94\epsilon\sqrt{\alpha}}{L\sqrt{2}},$ the distribution $\nu_{k}$ of the kth iterate of the ULMC algorithm satisfies $W_{2}^{2}(\nu_{k},\pi)\leq\frac{\epsilon^{2}d}{\alpha}$ for $k\geq c_{3}\frac{\gamma}{\alpha h}\log\frac{\sqrt{2\alpha}W_{2}(\nu_{0},\pi)}{\epsilon\sqrt{d}}$ and some absolute constant $c_{3}.$
With $K$ defined as in the corollary we have $\frac{\eta}{K}=\frac{0.94\epsilon\sqrt{\alpha}}{L\sqrt{2}},$ thus $W_{2}^{2}(U_{NK},\pi)\leq\frac{\epsilon^{2}d}{\alpha}$ for $N\geq\frac{c_{3}\gamma}{\alpha h}\log\frac{\sqrt{2\alpha}W_{2}(\nu_{0},\pi)}{\epsilon\sqrt{d}}.$ It remains to be shown that $W_{2}^{2}(\tilde{X}_{NK},X_{NK})\lesssim\frac{\epsilon^{2}d}{\alpha}.$ Let $k$ be a natural number. Under our assumptions on $V_{0}$ and $X_{0}$, we have $\mathbb{E}||V_{0}||^{2k}=d^{k}$ and
$(F(\Psi_{0})-F(\Psi_{T}))^{+}$ $\leq F(\Psi_{0})-F(x^{*})$
$\leq L||\Psi_{0}-x^{*}||^{2}$
$\lesssim L||X_{0}-x^{*}||^{2}+\frac{L}{\gamma^{2}}||V_{0}||^{2}$
Under our assumptions, we thus get $\mathbb{E}|(F(\Psi_{0})-F(\Psi_{T}))^{+}|^{k}\lesssim d^{k}.$ Moreover, we have $\log K=\log\frac{\eta L}{\epsilon\sqrt{\alpha}}\lesssim\log\frac{\kappa}{\epsilon}$ under the assumption that $\gamma\eta<c_{0}.$ Now let $L_{2}=c_{2}\log\frac{\sqrt{2\alpha}W_{2}(\nu_{0},\pi)}{\epsilon\sqrt{d}},$ $L_{3}=\log\frac{\kappa}{\epsilon}$ and apply Theorem 3 with $N$ as above, and
$\eta$ $\leq\min\Big{(}\tfrac{\epsilon^{1/3}}{\sqrt{L}},\tfrac{\epsilon^{1/3}}{\kappa^{1/6}L_{2}^{1/6}\sqrt{L}},\tfrac{\epsilon^{1/3}\kappa^{1/6}}{d^{1/6}\sqrt{L}},\tfrac{\epsilon^{1/3}}{\kappa^{1/6}d^{1/6}L_{2}^{1/3}\sqrt{L}},\tfrac{\epsilon^{\frac{p+2}{4p+3}}}{\kappa^{\tfrac{p/2-1}{4p+3}}d^{\tfrac{p}{4p+3}}\sqrt{L}}$
$\tfrac{\epsilon^{\tfrac{p+2}{4p+3}}}{\kappa^{\tfrac{3p}{8p+6}}d^{\tfrac{p}{4p+3}}L_{2}^{\tfrac{p+1}{4p+3}}\sqrt{L}},\tfrac{\epsilon^{1/3}d^{1/6}\kappa^{1/6}}{L_{3}^{1/6}\sqrt{L}},\tfrac{\epsilon^{1/3}d^{1/6}}{\kappa^{1/6}L_{3}^{1/3}\sqrt{L}},\tfrac{\epsilon^{\tfrac{p+2}{4p+3}}d^{\tfrac{1}{4p+3}}}{\kappa^{\tfrac{3p}{8p+6}}L_{2}^{\tfrac{p+1}{4p+3}}L_{3}^{\tfrac{p+1}{4p+3}}\sqrt{L}}\Big{)}.$
This gives $W_{2}^{2}(\tilde{X}_{NK},X_{NK})\leq\frac{\epsilon^{2}d}{\alpha},$ with $N=L_{3}\gamma(\alpha\eta)^{-1}$ as desired. ∎
Appendix G Deferred Technical Proofs
G.1 Proof of Proposition 2
The following proposition provides useful bounds on the operator norms of $\Gamma^{\prime}$ and $G^{\prime}$ based on Taylor series expansion. We refer to Section H for its proof.
Proposition 4.
[…]
  1. 1.
    $\|\Gamma^{\prime}_{h}\|^{2}\lesssim\frac{h}{\gamma}.$
  2. 2.
    $\|G_{h}^{\prime T}\Gamma_{h}^{\prime-2}G_{h}^{\prime}\|\leq\frac{h}{\gamma}.$
  3. 3.
    $\|G_{\gamma}^{\prime}[b(p)-b(q)]\|_{\mathbb{R}^{2d}}\lesssim\frac{hL}{\gamma}\|p_{u}-q_{u}\|_{\mathbb{R}^{d}}.$
We now prove Proposition 2
Proof.
Let $h=\frac{\eta}{K}$, and $T(X)=A_{h}^{\prime}X+G_{h}^{\prime}b(X),$ where $A_{h}^{\prime},G_{h}^{\prime}$ and $b$ are defined as above. Given Assumption 1, it is proven in [31] (Lemma 16) that the map $T$ is Lipschitz with $\|T\|_{\text{Lip}}\leq 1-\frac{\alpha}{\sqrt{L}}h+O(Lh^{2})$. Under our assumptions on $\eta$ and $\epsilon$ we have $L(\frac{\eta}{K})^{2}\lesssim\frac{\alpha}{\sqrt{L}}\cdot\frac{\eta}{K},$ and $T$ is thus a contraction with parameter $1-\Omega(\frac{\alpha\eta}{\sqrt{L}K}).$ Under the event $\mathcal{G}$, we have
$\|X_{tK+i+1}^{\prime}-\tilde{X}_{tK+i+1}^{\prime}\|^{2}$ $=\|T(x)-T(y)\|^{2}+\|\Gamma_{\frac{\eta}{K}}(Z_{tK+i}^{\prime}-\tilde{Z}_{tK+i}^{\prime})\|^{2}$
$+2\langle\Gamma_{\frac{\eta}{K}}^{\prime}(Z_{tK+i}^{\prime}-\tilde{Z}_{tK+i}^{\prime}),T(x)-T(y)\rangle.$
By the definition of $\tilde{Z}_{tK+i}$,
$\Gamma_{\frac{\eta}{K}}^{\prime}(\tilde{Z}_{tK+i}-Z_{tK+i})=\Gamma_{\frac{\eta}{K}}^{\prime}(Y_{tK+i}+V-Z_{tK+i})+G_{\frac{\eta}{K}}^{\prime}[b(\hat{\tilde{X}}_{tK+i})-b({\tilde{X}}_{tK+i})].$
Conditioned on $\mathcal{G}$, $(H_{i,t}-1/K)$ is zero mean and $Y_{tK+i}$ and $Z_{tK+i}$ are standard Gaussians. This gives $\mathbb{E}[\Gamma_{\frac{\eta}{K}}^{\prime}(Z_{tK+i}^{\prime}-\tilde{Z}_{tK+i}^{\prime})|\mathcal{G}]=G_{\frac{\eta}{K}}^{\prime}[b(\hat{\tilde{X}}_{tK+i}^{\prime})-b({\tilde{X}}_{tK+i}^{\prime})]$. By item 3 of Proposition 4, and the contractivity of $T$, we get
$\mathbb{E}[\|X_{tK+i+1}^{\prime}-\tilde{X}_{tK+i+1}^{\prime}\|^{2}|\mathcal{G}]$ $\leq(1-\Omega(\frac{\alpha\eta}{\gamma K}))\|x-y\|^{2}+O\Big{[}\|\Gamma_{\frac{\eta}{K}}(Z_{tK+i}^{\prime}-\tilde{Z}_{tK+i}^{\prime})\|^{2}$
$+\frac{\eta L}{K\gamma}\|\hat{u}-\tilde{u}\|\cdot\|x-y\|\Big{]}.$
An application of the AM-GM inequality gives
$\frac{\eta L}{\gamma K}\|\hat{u}-\tilde{u}\|\cdot\|x-y\|\lesssim\frac{\eta L^{2}}{\tau\alpha\gamma K}\|\hat{u}-\tilde{u}\|^{2}+\tau\frac{\alpha\eta}{\gamma K}\|x-y\|^{2},$
Where $\tau>0$ is arbitrary. Choose $\tau$ small enough so that the second term can be absorbed into $(1-\Omega(\frac{\alpha\eta}{\gamma K}))\|x-y\|^{2}$. We also have
$\mathbb{E}[\|\Gamma_{\frac{\eta}{K}}^{\prime}(\tilde{Z}_{tK+i}-Z_{tK+i})\|^{2}|\mathcal{G}]$ $\lesssim\|G_{\frac{\eta}{K}}^{\prime}[b(\hat{\tilde{X}}_{tK+i})-b({\tilde{X}}_{tK+i})]\|^{2}$
$+\|\Gamma_{\frac{\eta}{K}}^{\prime}\|^{2}\cdot\mathbb{E}[\|Y_{tK+i}+V-Z_{tK+i}\|^{2}|\mathcal{G}].$
By item 3 of Proposition 4, $\|G_{\frac{\eta}{K}}^{\prime}[b(\hat{\tilde{X}}_{tK+i})-b({\tilde{X}}_{tK+i})]\|^{2}\leq\frac{\eta^{2}L^{2}}{K^{2}\gamma^{2}}\|\tilde{u}-\hat{u}\|^{2}\leq\frac{\eta L^{2}}{\alpha\gamma K}\|\tilde{u}-\hat{u}\|^{2};$ and by item 1, $\|\Gamma_{\frac{\eta}{K}}^{\prime}\|^{2}\leq\frac{\eta}{\gamma K}.$
$\mathbb{E}[\|\Gamma_{\frac{\eta}{K}}^{\prime}(\tilde{Z}_{tK+i}-Z_{tK+i})\|^{2}|\mathcal{G}]$ $\lesssim\frac{\eta L^{2}}{\alpha\gamma K}\|\tilde{u}-\hat{u}\|^{2}+\frac{\eta}{\gamma K}\mathbb{E}[\|Y_{tK+i}+V-Z_{tK+i}\|^{2}|\mathcal{G}].$
Appendix H Proof of Proposition 4
Proof.
The eigenvalues of $\Gamma_{h}^{\prime 2}$ are
$E_{1}=\frac{\exp(-2\gamma h)}{\gamma^{2}}(a-b),$ $E_{2}=\frac{\exp(-2\gamma h)}{\gamma^{2}}(a+b)$
where $a=-1+\exp(2\gamma h)(1+2\gamma h),$ and
$b=\sqrt{1-32\exp(3\gamma h)+2\exp(2\gamma h)(7+2\gamma h)+\exp(4\gamma h)(17-4\gamma h+4\gamma^{2}h^{2})}.$
Taylor expansion in the variable $h$ gives
$\frac{\exp(-2\gamma h)}{\gamma^{2}}a=\frac{4h}{\gamma}-2h^{2}+\frac{4\gamma h^{3}}{3}+O(\gamma^{2}h^{4}),$
$\frac{\exp(-2\gamma h)}{\gamma^{2}}b=\frac{4h}{\gamma}-2h^{2}+\frac{7\gamma h^{3}}{6}+O(\gamma^{2}h^{4}).$
As a result, the eigenvalues $E_{1}$ and $E_{2}$ are of order $\gamma h^{3}$ and $\frac{h}{\gamma}$ respectively, with $E_{2}\geq E_{1}$ being the spectral norm of $\Gamma_{h}^{\prime 2}.$ We compute the inverse:
$\det(\Gamma_{h}^{\prime 2})=\frac{8}{\gamma^{4}}\exp(-2\gamma h)(-1+\exp(\gamma h))(2+\gamma h+\exp(\gamma h)(-2+\gamma h)),$
$\Gamma_{h}^{\prime-2}=\det(\Gamma_{h}^{\prime 2})^{-1}\begin{bmatrix}\frac{4(1-\exp(-\gamma h)-(1-\exp(2\gamma h))+2\gamma h}{\gamma^{2}}\mathbf{I}_{d}&-\frac{2\gamma h-(1-\exp(2\gamma h))}{\gamma^{2}}\mathbf{I}_{d}\\ -\frac{2\gamma h-(1-\exp(2\gamma h))}{\gamma^{2}}\mathbf{I}_{d}&\frac{4(1-\exp(-\gamma h)+(1-\exp(2\gamma h))+2\gamma h}{\gamma^{2}}\mathbf{I}_{d}\end{bmatrix}.$
An explicit computation gives
$G_{h}^{\prime T}\Gamma_{h}^{\prime-2}G_{h}^{\prime}=\begin{bmatrix}\frac{h}{2\gamma}&0\\ 0&0\end{bmatrix}.$
A Taylor expansion on the entries of $G_{h}^{\prime}$ shows
$G_{h}^{\prime}=\begin{bmatrix}\frac{h^{2}}{2}+O(\gamma h^{3})&0\\ \frac{2h}{\gamma}-\frac{h^{2}}{2}+O(\gamma h^{3})&0\\ \end{bmatrix}.$
Item 3 of the proposition follows from this and the smoothness of $F$ – Assumption 1. ∎